### v1

# Jet definition and TMD factorisation in SIDIS

Abstract
Using the colour dipole picture of Deep Inelastic Scattering (DIS) and the Colour Glass Condensate effective theory, we study semi-inclusive jet production in DIS at small $x$ in the limit where the photon virtuality $Q^{2}$ is much larger than the transverse momentum squared $P_{\perp}^{2}$ of the produced jet. In this limit, the cross-section is dominated by aligned jet configurations, that is, quark-antiquark pairs in which one of the fermions — the would-be struck quark in the Breit frame — carries most of the longitudinal momentum of the virtual photon. We show that physically meaningful jet definitions in DIS are such that the effective axis of the jet sourced by the struck quark is controlled by its virtuality rather than by its transverse momentum. For such jet definitions, we show that the next-to-leading order (NLO) cross-section admits factorisation in terms of the (sea) quark transverse momentum dependent (TMD) distribution, which in turn satisfies a universal DGLAP and Sudakov evolution.
Introduction.
The semi-inclusive production of a hadron or a jet in Deep Inelastic Scattering (SIDIS) is a process of fundamental importance for the study of the partonic content inside a proton or a large nucleus. It allows for an unambiguous extraction of the quark and gluon distribution functions with increasing accuracy in perturbative QCD Altarelli et al. (1979); Baier and Fey (1979); de Florian et al. (1998); Goyal et al. (2024); Bonino et al. (2024). In the small transverse momentum region of the measured hadron or jet, it is sensitive to the transverse momentum dependent (TMD) quark distribution  Collins (1993); Ji et al. (2005), whose precise extraction for both unpolarised and polarised target is one of the main goal of the Electron-Ion Collider (EIC) physics program Accardi et al. (2016); Abdul Khalek et al. (2022); Achenbach et al. (2024). In this Letter, we consider the SIDIS process in the case of a jet measurement at small $x_{\rm Bj}$, where the CGC effective theory Iancu et al. (2002); Iancu and Venugopalan (2003); Gelis et al. (2010); Kovchegov and Levin (2013) applies, and demonstrate that the factorisation of the NLO cross-section in terms of the (sea) quark TMD necessitates the use a new jet reconstruction algorithm which captures the non-trivial dynamics of the quark struck by the virtual photon.
Compared to hadron production in SIDIS, jet process has a significant advantage because it can directly probe the TMD quark distribution without involving a TMD fragmentation function. However, a precise measurement of low transverse momentum jets is very challenging, in particular because the usual $k_{t}$-type algorithms may not be applicable. Alternative approaches and detailed phenomenology applications have been carried out in the last few years Gutierrez-Reyes et al. (2018, 2019a, 2019b); Liu et al. (2019, 2020); Aschenauer et al. (2020); Page et al. (2020); Arratia et al. (2020a, 2021, b); Kang et al. (2021a, 2020, b); Andreev et al. (2022); Lai et al. (2022); Lee et al. (2023); Arratia et al. (2023); Fang et al. (2024). Together with these developments, the novel jet algorithm proposed in this paper will help to improve the precision of future jet measurements in SIDIS at the EIC and provide important probe to the nucleon/nucleus tomography in terms of the TMD quark distributions at various $x$ ranges.
Aligned jet configurations.
We work in the dipole frame (related to the Breit frame by a longitudinal boost) where the virtual photon has four-momentum $q^{\mu}=(q^{+},q^{-}=-Q^{2}/(2q^{+}),\bm{0}_{\perp})$ and a nucleon from the nucleus target has four momentum $P^{\mu}_{N}=(0,P^{-}_{N},\bm{0}_{\perp})$ in light-cone coordinates. The standard DIS variables are defined by $Q^{2}=-q^{2}$ and $x_{\rm Bj}=Q^{2}/\hat{s}$ with $\hat{s}=2q^{+}P^{-}_{N}$. At LO in the colour dipole picture of DIS at small $x_{\rm Bj}$ Kopeliovich et al. (1981); Bertsch et al. (1981); Mueller (1990); Nikolaev and Zakharov (1991), the transversely polarised virtual photon $\gamma_{\textrm{T}}^{*}$ splits into a quark-antiquark pair which interacts with the nucleus target. We denote by $k_{1,2}^{\mu}$ the four-momenta of the quark and the antiquark and define $z_{i}=k_{i}^{+}/q^{+}$. Each outgoing parton can then form a jet for which one only measures the transverse momentum $\bm{P}_{\perp}$. We are primarily interested in the limit $Q^{2}\gg P_{\perp}^{2}$. The leading power in the $P_{\perp}^{2}/Q^{2}$ expansion of the cross-section comes from aligned-jet configurations Mueller (1999); Marquet et al. (2009); Iancu et al. (2021), that is very asymmetric $q\bar{q}$ pairs such that the quark has $1-z_{1}\sim P_{\perp}^{2}/Q^{2}\ll 1$ and carries most of the longitudinal momentum of the virtual photon while the antiquark has $z_{2}\sim P_{\perp}^{2}/Q^{2}\ll 1$, or vice-versa (cf. Fig. 1). The struck fermion in the target picture is naturally the one with $z_{i}\sim 1$, yet the tagged jet with transverse momentum $\bm{P}_{\perp}$ can be either the fast ($z_{i}\sim 1$), or the slow ($z_{i}\ll 1$), fermion in the dipole picture. In the limit $Q^{2}\gg P_{\perp}^{2}$, the semi-inclusive single jet cross-section admits TMD factorisation Marquet et al. (2009) in terms of the sea quark TMD $x\mathcal{F}_{q}(x,\bm{P}_{\perp})$ at small $x$ McLerran and Venugopalan (1999); Venugopalan (1999); Mueller (1999):
$\left.\frac{\mathrm{d}\sigma^{\gamma^{*}_{\rm T}+A\to j+X}}{\mathrm{d}^{2}\bm{P}_{\perp}}\right|_{\rm LO}$ $=\frac{8\pi^{2}\alpha_{\rm em}e_{f}^{2}}{Q^{2}}x\mathcal{F}_{q}(x,\bm{P}_{\perp})\,,$ (1)
where $x=x_{\rm Bj}$ as a consequence of minus momentum conservation. $\alpha_{\rm em}$ is the fine structure constant and $e_{f}$ is the fractional electric charge of a light quark. This expression covers the case where the quark jet is measured  [^1] A similar expression holds for the case where the antiquark jet is measured. See Cacciari et al. (2019) for a discussion of the unitarity of the SIDIS cross-section with jets., but it is inclusive in the jet longitudinal momentum fraction $z$. Explicit expression for the sea quark TMD in term of the quark-antiquark dipole $S$-matrix at small $x$ can be found in Xiao et al. (2017); Marquet et al. (2009); Tong et al. (2023) (see also the Supplemental Material for a brief review).
Figure 1: Geometric representation of the typical aligned jet configuration at small $x_{\rm Bj}$ in the dipole frame.
Heuristic discussion of the Sudakov logarithm.
At NLO in the strong coupling $\alpha_{s}$, the dominant radiative corrections to Eq. (1) are enhanced by double Sudakov logarithms Sudakov (1956); Dokshitzer et al. (1980); Parisi and Petronzio (1979); Collins et al. (1985) in the ratio of $Q^{2}/P_{\perp}^{2}$. To DLA, these Sudakov logarithms come from virtual emissions by the struck quark in the phase-space that is forbidden to the real gluon emissions, that is real gluons which would modify the structure of the final state Mueller et al. (2013a, b). Clearly, the real gluon emissions with transverse momenta such that $k_{g\perp}^{2}\gg P_{\perp}^{2}$ must be forbidden. On the other hand, virtual emissions with $k_{g\perp}^{2}\geq Q^{2}$ do not have any double logarithmic support (the UV logarithmic divergence cancels above the hard scale $Q^{2}$ among virtual graphs Caucal et al. (2024)). The transverse phase-space which contributes to the Sudakov double logarithm is therefore $P_{\perp}^{2}\ll k_{g\perp}^{2}\ll Q^{2}$. Concerning the longitudinal phase space, we observe that the gluon emission must have a formation time $\tau_{g}=1/k_{g}^{-}$ larger than the coherence time $\tau_{\gamma}=1/|q^{-}|$ of the virtual photon Beuf (2014); Iancu et al. (2015); Hatta and Iancu (2016); Ducloué et al. (2019). In the context of high energy factorisation, this condition is imposed on any gluon emission which does not contribute to the (collinearly improved) BK/JIMWLK evolution Balitsky (1996); Kovchegov (1999); Jalilian-Marian et al. (1997, 1998); Kovner et al. (2000); Weigert (2002); Iancu et al. (2001a, b); Ferreiro et al. (2002) of the target wave-function. In terms of $z_{g}=k_{g}^{+}/q^{+}$, this gives the constraint $z_{g}\gg k_{g\perp}^{2}/Q^{2}$. We thus end up with the following integral for the Sudakov double logarithm,
$S_{\rm DL}=-\frac{\alpha_{s}C_{F}}{2\pi}\int_{P_{\perp}^{2}}^{Q^{2}}\frac{\mathrm{d}k_{g\perp}^{2}}{k_{g\perp}^{2}}\int_{k_{g\perp}^{2}/Q^{2}}^{z_{\rm max}}\frac{\mathrm{d}z_{g}}{z_{g}}\,,$ (2)
which factorises from the LO cross-section Eq. (1).
Consider now the upper limit $z_{\rm max}$ of the $z_{g}$ integral. In the case of a hadron measurement, a forbidden real gluon must be well separated from the collinear singularity for a final-state emission. This singularity corresponds to (soft) gluons obeying $\bm{k_{g\perp}}=z_{g}\bm{P}_{\perp}/z_{1}$, with $z_{1}\simeq 1$ for the struck quark. Clearly such collinear gluons have no overlap with the forbidden phase-space at $k_{g\perp}^{2}\gg P_{\perp}^{2}$, so the collinear singularity introduces no additional constraint on this phase-space. We can then take $z_{\rm max}=1$, which gives
$S_{\rm DL}^{\rm had}=-\frac{\alpha_{s}C_{F}}{2\pi}\ln^{2}\left(\frac{Q^{2}}{P_{\perp}^{2}}\right),$ (3)
as expected for hadron production Xiao et al. (2017); Altinoluk et al. (2024).
The case of a jet measurement is considerably more subtle, and here comes our main physical observation. The fast quark with $z_{1}\sim 1$ (see Fig. 1) is not put on-shell by the scattering, rather it emerges from the collision with a relatively large virtuality, of order $Q^{2}$. Indeed, after the photon decay $\gamma\to q\bar{q}$, the two quarks separate via quantum diffusion. By the time $\tau_{\gamma}=2q^{+}/Q^{2}$ of the collision with the target, the wave packets of the fast quark and the slow antiquark spread out over distances $\Delta x_{q\perp}^{2}\sim\tau_{\gamma}/(z_{1}q^{+})$ and $\Delta x_{\bar{q}\perp}^{2}\sim\tau_{\gamma}/(z_{2}q^{+})$, respectively. In the aligned jet limit, $\Delta x_{q\perp}^{2}\sim 1/Q^{2}$ and $\Delta x_{\bar{q}\perp}^{2}\sim 1/P_{\perp}^{2}$. For the fast quark, $\Delta x_{q\perp}^{2}\ll 1/P_{\perp}^{2}$ is very small, showing that this parton is still localised by its virtuality. Physically, this is so since the collision occurs relatively fast: $\tau_{\gamma}=2q^{+}/Q^{2}$ is much smaller than the formation time $\tau_{q}\sim 2z_{1}q^{+}/P_{\perp}^{2}$ for the fast quark. Accordingly, the angle made by this quark with respect to the collision axis by the time of scattering can be estimated as $\theta_{q}\sim\Delta x_{q\perp}/\tau_{\gamma}\sim Q/q^{+}$, which is much bigger than the naive angle $\theta_{0}\sim P_{\perp}/q^{+}$ it would have made if it was on-shell (i.e.  localised by its transverse momentum, see Fig. 1).
In order to be forbidden, a real gluon emission must not be part of the jet sourced by the fast quark. This condition imposes $\theta_{g}\sim k_{g\perp}/k_{g}^{+}\gg\theta_{q}$, i.e. $z_{g}\ll k_{g\perp}/Q$. Using $z_{\rm max}=k_{g\perp}/Q$ in Eq. (2), one finds
$S_{\rm DL}^{\rm jet}$ $=-\frac{\alpha_{s}C_{F}}{4\pi}\ln^{2}\left(\frac{Q^{2}}{P_{\perp}^{2}}\right)\,.$ (4)
Thus, remarkably, the Sudakov corresponding to a jet final state is smaller by a factor of 2 than that for a hadron final state. This is so since the fast virtual quark generates a relatively wide jet, so the phase-space that is forbidden to real emission is correspondingly smaller.
A new jet distance measure in DIS.
To properly reflect the above physical picture, a jet measurement in DIS must be endowed with a clustering algorithm which accounts for the high virtuality $\sim Q^{2}$ of the struck quark. This is generally not the case for the jet algorithms that are a priori designed for hadron-hadron collisions, such as the $k_{t}$-algorithms Ellis and Soper (1993); Dokshitzer et al. (1997); Wobisch and Wengler (1998); Cacciari et al. (2008, 2012): when applied to DIS, they typically cluster particles ”around” an axis with angle $P_{\perp}/(zq^{+})$ in the dipole frame Ivanov and Papa (2012); Boussarie et al. (2016); Roy and Venugopalan (2020a, b); Caucal et al. (2021); Taels et al. (2022); Caucal et al. (2022); Liu et al. (2022) and thus do not capture the effect of the large virtuality of the jet with $z\sim 1$. Related to that, they fail to cluster the remnant of the struck quark into the same jet, as noted in Arratia et al. (2021). In order to cope with these issues, we introduce a new jet distance measure, via
$d_{ij}=\frac{M_{ij}^{2}}{(z_{i}z_{j})^{p}Q^{2}R^{2}}\,,\quad d_{iB}=1\,,$ (5)
where $d_{ij}$ is the distance between two particles labelled $i,j$, and $d_{iB}$ is the particle-beam distance  [^2] The measure corresponds to the particular case $k=0$ of a generalised distance measure $d_{ij}=\textrm{min}(k_{i\perp}^{2k},k_{j\perp}^{2k})M_{ij}^{2}/((z_{i}z_{j})^{p}Q^{2}R^{2})$, $d_{iB}=k_{i\perp}^{2k}$ as done in hadron-hadron collisions Cacciari et al. (2008); Salam (2010).. The measure depends on the invariant mass squared $M_{ij}^{2}=(k_{i}+k_{j})^{2}$ and the longitudinal momentum fractions $z_{i}=k_{i}\cdot P_{N}/(q\cdot P_{N})$. For a given list of final state particles, the jet algorithm then runs inclusively Salam (2010) through pairwise recombination of particles $i_{0},j_{0}$ if $d_{i_{0}j_{0}}$ is minimal among all $d_{ij}$,$d_{iB}$, $d_{jB}$, while declaring $i_{0}$ a final state jet if $d_{i_{0}B}$ is minimal, until no particle remains.
The algorithm depends on two parameters, the jet radius $R$ and the power $p$. With $p=-1,0,1$, it gives respectively a $k_{t}$, mass, or angular ordered clustering in the dipole frame. We shall focus here on the specific choice $p=1$, which is dynamically favoured, as we shall shortly argue. With this choice, one finds (in the dipole frame and in the limit of a small relative angle $\theta_{ij}\ll 1$)
$d_{ij}\approx\frac{\theta_{ij}^{2}(q^{+})^{2}}{Q^{2}R^{2}}\,,\quad\textrm{ for }p=1\,.$ (6)
Using this criterion at NLO, where the final state is made of two partons, the quark $i$ and gluon $j$, one sees that these two partons are clustered within the same jet if $\theta_{ij}\leq R\theta_{q}$ with $\theta_{q}\sim Q/q^{+}$, in agreement with the dynamics of the fast virtual quark that we have just elucidated. This is a consequence of both the normalisation of the distance measure with $Q^{2}$ and the choice of $p$. We assume the jet radius $R$ to be small, but not too small, so that an additional $\ln(1/R)$ resummation is not needed Dasgupta et al. (2015)  [^3] For jet production with a very small-$R$, one needs to introduce the so-called semi-inclusive jet function Kang et al. (2016) and the factorisation will be different from what we discuss in this paper..
For $p=1$ and in the fragmentation region of the struck quark in the Breit frame, the algorithm is akin to generalised $k_{t}$ algorithms originally designed for $e^{+}e^{-}$ annihilation Bartel et al. (1986); Bethke et al. (1988); Catani et al. (1991); Cacciari et al. (2012) and subsequently extended to DIS in Catani et al. (1992); Webber (1993) (cf Supplemental Material). However, unlike the latter, which are only defined in the Breit frame, Eq. (5) is longitudinally invariant by boost along the $\gamma^{*}$-$A$ collision axis (see also Arratia et al. (2021)), so that $d_{ij},d_{iB}$ can be computed either in the Breit frame, or the dipole frame. This makes the distance measure Eq. (5) more convenient for higher order calculations and jet studies in DIS.
TMD factorisation at NLO.
For the inclusive production of jets in DIS, that is, in situations where the jet transverse momenta are not measured, but only their number, it was known since Catani et al. (1992); Webber (1993) that special jet definitions are needed to guarantee the validity of the standard collinear factorisation for the jet cross-section. In what follows, we would like to demonstrate a similar property for the semi-inclusive production of a jet with a given $P_{\perp}$. Specifically, we will show that the NLO corrections to the jet cross-section (as computed in the dipole picture) are consistent with TMD factorisation — in the sense that they correctly generate the expected DGLAP Gribov and Lipatov (1972); Altarelli and Parisi (1977); Dokshitzer (1977) and CSS Collins and Soper (1981, 1982); Collins et al. (1985); Collins (2023) evolutions of the quark TMD — if and only if the quark and the gluon jets are separated from each other by using the jet distance (5) with $p=1$. To show that, we consider a relatively hard jet, with $Q^{2}\gg P_{\perp}^{2}\gg Q_{s}^{2}$. In this regime and to lowest order in $Q_{s}^{2}/P_{\perp}^{2}$, the transverse momentum of the measured jet is either given by the recoil from a hard gluon emission with $\bm{k_{g\perp}}\sim-\bm{P}_{\perp}$ (real contribution) or by the target itself (virtual contribution).
For the real contribution, our starting point is the quark-antiquark-gluon Fock component of the virtual photon wavefunction and, more precisely, its contribution to the $\gamma_{\rm T}^{*}+A\to qg$ cross-section, as obtained after integrating out the antiquark  [^4] If instead of the antiquark, one integrates out the gluon, that is, if one computes a real-gluon correction to the $\gamma_{\rm T}^{*}+A\to qq$ cross-section, then one generates a step in the DGLAP+CSS evolution of the gluon TMD Caucal and Iancu (2024).. The quark and the gluon are nearly back-to-back: their transverse momentum imbalance $\bm{\ell}_{\perp}=\bm{k_{g\perp}}+\bm{k_{1\perp}}$ is small compared to the relative momentum $\bm{P}_{\perp}=z_{g}\bm{k_{1\perp}}-z_{1}\bm{k_{g\perp}}$: $\ell_{\perp}\ll P_{\perp}$ and $z_{1}+z_{g}\simeq 1$. In the CGC calculation, this process is shown to factorise in terms of the quark TMD as Hauksson et al. (2024); P. Caucal, E. Iancu, F. Salazar, and F. Yuan (2024)
$\frac{\mathrm{d}\sigma^{\gamma_{\textrm{T}}^{\star}+A\to qg+X}}{\mathrm{d}^{2}\bm{P}_{\perp}\mathrm{d}^{2}\bm{\ell}_{\perp}\mathrm{d}z_{1}\mathrm{d}z_{g}}=\alpha_{\rm em}e_{f}^{2}\alpha_{s}C_{F}\delta(1-z_{1}-z_{g})$
×$\times\frac{2z_{1}\left[(P_{\perp}^{2}+\bar{Q}^{2})^{2}+z_{g}^{2}P_{\perp}^{4}+z_{1}^{2}\bar{Q}^{4}\right]}{P_{\perp}^{2}\left[P_{\perp}^{2}+\bar{Q}^{2}\right]^{3}}x_{q}\mathcal{F}_{q}(x_{q},\bm{\ell}_{\perp})\,,$ (7)
with $\bar{Q}^{2}=z_{1}z_{g}Q^{2}$ and $x_{q}=(M_{qg}^{2}+Q^{2})/\hat{s}$. The contribution of (7) to SIDIS is obtained by integrating out the kinematics of the unmeasured jet over the region in phase-space where the quark and the gluon form two separated jets. Using the jet definition Eq. (5) with $p=1$, the associated phase space constraint reads $P_{\perp}^{2}\geq R^{2}Q^{2}z_{1}^{2}z_{g}^{2}$. The subsequent integrals over $z_{1}$ and $z_{g}$ are controlled by $z_{1}\simeq 1$ and $z_{g}\lesssim P_{\perp}/(RQ)\ll 1$, so that only very soft gluons contribute. Such soft gluons can be transferred into the target wavefunction Iancu et al. (2022a, b); Hauksson et al. (2024); Caucal and Iancu (2024) via a change of variable $z_{g}\to\xi$, with
$z_{g}=\frac{\xi}{1-\xi}\frac{P_{\perp}^{2}}{Q^{2}}\ \ \Longleftrightarrow\ \ \xi=\frac{x_{\rm Bj}}{x_{q}}$ (8)
where we have also used $M_{qg}^{2}=P_{\perp}^{2}/(z_{1}z_{g})$ and $z_{1}\simeq 1$. The integral over $\xi$ is restricted to $x_{\star}<\xi<1-RP_{\perp}/Q$, where the upper limit is introduced by the jet constraint $z_{g}\lesssim P_{\perp}/(RQ)$, while the lower limit comes from the separation of the longitudinal phase-space between the NLO impact factor and the high energy evolution. This separation is more precisely written as $z_{g}\geq x_{\star}P_{\perp}^{2}/Q^{2}$, with $x_{\star}$ an arbitrary factorisation scale which obeys $x_{\star}\ll 1$ and $\alpha_{s}\ln(1/x_{\star})\ll 1$. Finally, the integral over the momentum imbalance $\ell_{\perp}\ll P_{\perp}$ builds the quark PDF $xf_{q}$ on the resolution scale of the hard jet:
$xf_{q}(x,P_{\perp}^{2})=\int_{\Lambda^{2}}^{P_{\perp}^{2}}{\mathrm{d}^{2}\bm{\ell}_{\perp}}\,x\mathcal{F}_{q}(x,\bm{\ell}_{\perp})\,,$ (9)
with $\Lambda$ of the order of the QCD confinement scale. Based on the above, a straightforward calculation yields
$\left.\frac{\mathrm{d}\sigma^{\gamma^{*}_{\rm T}+A\to j+X}}{\mathrm{d}^{2}\bm{P}_{\perp}}\right|_{R}=\frac{8\pi^{2}\alpha_{\rm em}e_{f}^{2}}{Q^{2}}$
×$\times\frac{\alpha_{s}}{2\pi^{2}}\frac{1}{P_{\perp}^{2}}\int_{x_{\star}}^{1-\frac{RP_{\perp}}{Q}}{\mathrm{d}\xi}P_{qq}(\xi)\frac{x_{\rm Bj}}{\xi}f_{q}\left(\frac{x_{\rm Bj}}{\xi},P_{\perp}^{2}\right)\,,$ (10)
with $P_{qq}(\xi)=C_{F}(1+\xi^{2})/(1-\xi)$ the unregularised $q\to qg$ splitting function. In Eq. (10) we included a factor of two to account for the fact that the tagged jet can be generated by either the fast quark ($z_{1}\simeq 1$), or the slow gluon ($z_{g}\ll 1$). The NLO correction in Eq. (10) exhibits TMD factorisation, as anticipated: the expression in the second line can be interpreted as an evolution of the quark TMD. The singularity of $P_{qq}(\xi)$ at $\xi=1$ introduces a logarithmic sensitivity to the upper limit, that ca be isolated with the help of the plus prescription:
$\left.x\mathcal{F}_{q}^{(1)}(x,\bm{P}_{\perp},Q^{2})\right|_{R}=\frac{\alpha_{s}}{2\pi^{2}}\frac{1}{P_{\perp}^{2}}\!\int_{x_{\star}}^{1}{\mathrm{d}\xi}P_{qq}^{(+)}(\xi)\frac{x}{\xi}f_{q}\left(\frac{x}{\xi},P_{\perp}^{2}\right)$
$\quad\qquad\qquad+\frac{\alpha_{s}C_{F}}{2\pi^{2}}\frac{1}{P_{\perp}^{2}}\ln\left(\frac{Q^{2}}{R^{2}P_{\perp}^{2}}\right)xf_{q}(x,P_{\perp}^{2})\,,$ (11)
where $P_{qq}^{(+)}(\xi)$ differs from $P_{qq}(\xi)$ only via the replacement $(1-\xi)\to(1-\xi)_{+}$ in the denominator. We recognise in Eq. (11) one (real) step in the DGLAP+CSS evolution of the quark TMD. (See also Hauksson et al. (2024); Caucal and Iancu (2024) for similar arguments.) The $1/P_{\perp}^{2}$ behaviour comes from the hard $q\to qg$ splitting leading to two well separated jets. The precise coefficient of the Sudakov logarithm in the second line is a consequence of choosing $p=1$ in our jet distance measure (5). If one were using another jet definition from that class, or a definition with an effective jet axis set by the angle $\theta_{0}$ in Fig. 1, the upper limit of the $\xi$ integral in Eq. (10) would change (e.g., it would be $\xi<1-R^{2}P_{\perp}^{2}/Q^{2}$ for $p=0$), which would in turn modify the normalisation of the Sudakov logarithm. With $p=-1$ in Eq. (5), the quark and the gluon would typically be clustered within the same jet, so that the TMD evolution that we have just unveiled would not be resolved within the wide jet formed by the struck quark.
An important consistency check refers to the detailed balance between the real and the virtual NLO corrections: after integrating the cross-section over $\bm{P}_{\perp}$, the Sudakov logarithms must cancel between real and virtual terms (see e.g. the discussion in Caucal and Iancu (2024)). In order to verify this condition and also to complete our calculation of the cross-section to the accuracy of interest, we need the virtual NLO contributions in the limit $Q^{2}\gg P_{\perp}^{2}\gg Q_{s}^{2}$. They can be inferred from the NLO calculation of the SIDIS cross-section in the CGC, as presented in Caucal et al. (2024). The details of the leading power extraction in these expressions are provided in Supplemental Material. The virtual terms too are sensitive to the jet definition, as clear from the fact that they include the phase-space region where the quark and the gluon are nearly collinear with each other and hence must be clustered within the same jet. The collinear singularity cancels between real and virtual contributions, but the finite reminder, which is enhanced by (double and single) Sudakov logarithms, is clearly dependent upon our definition for the jets. Using Eq. (5) with $p=1$, the virtual term reads
$\left.\frac{\mathrm{d}\sigma^{\gamma^{*}_{\rm T}+A\to j+X}}{\mathrm{d}^{2}\bm{P}_{\perp}}\right|_{V}=\left.\frac{\mathrm{d}\sigma^{\gamma^{*}_{\rm T}+A\to j+X}}{\mathrm{d}^{2}\bm{P}_{\perp}}\right|_{\rm LO}$$\times\frac{\alpha_{s}C_{F}}{\pi}\left[-\frac{1}{4}\ln^{2}\left(\frac{Q^{2}}{P_{\perp}^{2}}\right)+\left(\frac{3}{4}+\ln(R)\right)\ln\left(\frac{Q^{2}}{P_{\perp}^{2}}\right)\right.$
$\left.-\frac{3}{2}\ln(R)+\frac{11}{4}-\frac{3\pi^{2}}{4}+\frac{3}{4}\ln^{2}(x_{\star})+\frac{3}{8}\ln(x_{\star})+\mathcal{O}(R^{2})\right]\,.$ (12)
The first term in the square bracket in Eq. (12) is the Sudakov double logarithm that we heuristically derived in Eq. (4). It agrees with the expectation from the CSS kernel for the quark TMD alone Collins and Soper (1981, 1982); Collins et al. (1985); Collins (2023). The second term is a Sudakov single logarithm, which depends both on the quark anomalous dimension $\Gamma_{q}=\frac{3\alpha_{s}C_{F}}{4\pi}$ and on the jet parameter $R$. By only keeping these logarithmic terms, one obtains the virtual contribution to the DGLAP+CSS evolution of the quark TMD, conveniently rewritten as
$\left.x\mathcal{F}_{q}^{(1)}(x,\bm{P}_{\perp},Q^{2})\right|_{V}=-\frac{\alpha_{s}C_{F}}{2\pi}\,x\mathcal{F}_{q}(x,\bm{P}_{\perp})$
×$\quad\qquad\qquad\qquad\times\int_{P_{\perp}^{2}}^{Q^{2}}\frac{\mathrm{d}\ell_{\perp}^{2}}{\ell_{\perp}^{2}}\left[\ln\left(\frac{Q^{2}}{R^{2}P_{\perp}^{2}}\right)-\frac{3}{2}\right],$ (13)
After integrating Eqs. (11) and (13) over $P_{\perp}^{2}$ up to $Q^{2}$, it is straightforward to verify that the effects of the Sudakov logarithms which are manifest in these two equations mutually cancel, as announced. Furthermore, the piece $3/2$ in Eq. (13) gives a contribution $(3/2)\delta(1-\xi)$ to the splitting function in Eq. (11), thus completing the standard expression for the regularised DGLAP splitting function $\mathcal{P}_{qq}(\xi)$ Collins (2023). This real vs. virtual cancellation is a crucial feature of the jet definition that we employed, i.e. Eq. (5) with $p=1$ (we show in Supplemental Material that other common choices of DIS jet definitions do not satisfy this condition). We thus obtain the DGLAP equation for the quark PDF in integral form:
$xf_{q}(x,Q^{2})=$ $xf_{q}^{(0)}(x,Q^{2})+\int_{\Lambda^{2}}^{Q^{2}}\!\frac{\mathrm{d}P_{\perp}^{2}}{P_{\perp}^{2}}\frac{\alpha_{s}(P_{\perp}^{2})}{2\pi}$
×$\times\!\int_{x_{\star}}^{1}{\mathrm{d}\xi}\,\mathcal{P}_{qq}(\xi)\,\frac{x}{\xi}f_{q}\left(\frac{x}{\xi},P_{\perp}^{2}\right),$ (14)
where we have also inserted a running coupling, as standard in this context. The CSS equation obeyed by the quark TMD can be simply obtained by taking a derivative w.r.t. $\ln Q^{2}$ in the sum of Eqs. (11) and (13).
To summarise, we have demonstrated that TMD factorisation for jet production in SIDIS is not guaranteed by all jet definitions and we have designed a longitudinally-invariant jet clustering algorithm which preserves both the factorisation and the universality of the DGLAP+CSS evolution of the quark TMD at small $x$. Physically, this jet definition is able to resolve the DGLAP and Sudakov dynamics of the struck sea quark. While these results are derived in the context of the high energy factorisation, they also apply at moderate values of $x$, as demonstrated in the Supplemental Material, where we find a similar factorisation property by following the dynamics of the struck quark in the target picture. Once again, the proper choice of the jet definition turns out to be essential to ensure the correct matching between the NLO corrections to SIDIS and the DGLAP+CSS evolution of the quark TMD. It will be important to investigate further the jet definition Eq. (5) both at small and moderate $x$, as we anticipate that it could be of great importance for the forthcoming quark tomography at the EIC with jets Arratia et al. (2020a); Page et al. (2020).
Acknowledgements.
Acknowledgements. We are grateful to Cyrille Marquet, Farid Salazar, Gregory Soyez, and Werner Vogelsang for inspiring discussions. We thank the France-Berkeley-Fund from University of California at Berkeley for support. The work of F.Y. is supported in part by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, under contract number DE-AC02-05CH11231. The work of A.H.M. is supported in part by the U.S. Department of Energy Grant # DE-FG02-92ER40699.
Appendix A Supplemental material
In this Supplemental Material, we will discuss in more detail two topics that have been briefly addressed in the text of the Letter and which, as explained there, are deeply related to each other: the definition of jets in DIS and the TMD factorisation of the cross-section for SIDIS as computed in the colour dipole picture to next-to-leading order (NLO). We will notably show that some of the standard jet definitions used in the literature are inconsistent with TMD factorisation at NLO. In practice, this means that the respective definitions typically fail to properly reconstruct the jet structure generated by the struck quark.
Appendix B 1 Jet definitions and NLO clustering conditions
In this section, we briefly review various standard jet definitions which have been used in phenomenological studies of DIS Newman and Wing (2014); Chekanov et al. (2002, 2007); Abramowicz et al. (2010); Aktas et al. (2007); Andreev et al. (2017); Arratia et al. (2020a); Page et al. (2020); Abt et al. (2023); Arratia et al. (2021), with the purpose of comparing them with the new definition that we introduced in Eq. (5) of the Letter. In the original literature, these definitions are either formulated in the Breit frame, or, in some rare occasions, they are given Lorentz-invariant formulations. In what follows, we shall re-express them in the dipole frame, which will be convenient for the NLO calculations to be presented in the next section. We always consider massless particles for simplicity.
B.1 1.1 Longitudinally invariant jet algorithms in hadron-hadron collisions
Longitudinally invariant generalised-$k_{t}$ algorithms Ellis and Soper (1993); Dokshitzer et al. (1997); Wobisch and Wengler (1998); Cacciari et al. (2008, 2012) in inclusive mode  [^5] We consider throughout this section the inclusive mode Salam (2010) where a jet $i$ such that $d_{iB}$ is minimal is declared as a final jet, which is then removed from the list of particles. for hadron-hadron collisions are defined with the distance measure
$d_{ij}=\textrm{min}(k_{i\perp}^{2k},k_{j\perp}^{2k})\frac{\Delta R_{ij}^{2}}{R^{2}}\,,\quad d_{iB}=k_{i\perp}^{2k}$ (15)
where $\Delta R_{ij}^{2}$ is the square distance between the two particles $i$ and $j$ in the rapidity-azimuth plane. When there are only two particles in the final state, as it is the case at NLO for the SIDIS process we are considering, two particles $i$ and $j$ are recombined within a same jet if $d_{ij}\leq\textrm{min}(d_{iB},d_{jB})$, or, equivalently, $\Delta R_{ij}^{2}\leq R^{2}$. (In this situation, the power $k$ plays no role in the clustering condition, but one should keep in mind that when the final state is made of more than two particles, this power modifies the ordering variable of the clustering sequence  [^6] If there are more than 2 particles in the final state, the clustering condition is more complicated because $d_{ij}$ must also be smaller than all the others $d_{i^{\prime}j^{\prime}}$ for all pairs..) This algorithm is longitudinally invariant, so it can be implemented either in the Breit frame or in the dipole frame — it gives the same jets in both cases. For small $\Delta R_{ij}$, we have Liu and Petriello (2013); Liu et al. (2022)
$\Delta R_{ij}^{2}\approx\frac{2k_{i}\cdot k_{j}}{k_{i\perp}k_{j\perp}}$ (16)
so that the clustering condition $\Delta R_{ij}^{2}\leq R^{2}$ when only two particles are produced in the final state is
[…]
where $z$ and $\bm{P}_{\perp}$ are the jet longitudinal momentum fraction and transverse momentum in the so-called E-recombination scheme Catani et al. (1993); Blazey et al. (2000), in which the jet four-momentum is given by the sum of the four-momenta of the two clustered particles. To go from Eq. (16) to Eq. (17), we have used $M_{ij}^{2}\equiv(k_{i}+k_{j})^{2}=2k_{i}\cdot k_{j}$ for on-shell massless particles together with the collinear approximation where $k_{i\perp}\sim(z_{i}/z)P_{\perp}$. We shall refer to Eq. (17) as the clustering condition A.
B.2 1.2 Jet algorithms in DIS
B.2.1 1.2.1 JADE-like algorithm from Webber (1993)
A JADE-like algorithm Bartel et al. (1986); Bethke et al. (1988) for DIS has been proposed by Webber in Webber (1993). It relies on the measure
$d_{ij}=\frac{M_{ij}^{2}}{Q^{2}R^{2}}\,,\quad d_{iB}=2x_{\rm Bj}\frac{k_{i}\cdot P}{Q^{2}}\,,$ (18)
where the jet resolution parameter $d_{\rm cut}$ is written as $d_{\rm cut}=R^{2}$ following Webber (1993). One advantage of this JADE-like definition is that the distance measure is Lorentz-invariant. At NLO, two particles $i$ and $j$ are recombined if $d_{ij}\leq\textrm{min}(d_{iB},d_{jB})$. In the dipole frame, where $k_{i}\cdot P=z_{i}(q\cdot P)$ and hence $d_{iB}=z_{i}$, this condition amounts to
$\frac{M_{ij}^{2}}{Q^{2}R^{2}}\leq\textrm{min}(z_{i},z_{j})$ (19)
We shall not consider this clustering condition in our NLO calculation, as the function $\textrm{min}(z_{i},z_{j})$ in the right hand side makes the computation more cumbersome, but we have checked that it gives the same NLO results modulo pure $\mathcal{O}(\alpha_{s})$ corrections. In that sense, it is equivalent (at least to up to the NLO) to the jet definition with $p=1$ proposed in Eq. (5) in the main text. Note that the special choice $d_{iB}=z_{i}$ for the beam distance is crucial in that sense: if it were fixed to $1$ as in the JADE algorithm for $e^{+}e^{-}$ annihilation, the distance measure would be identical to Eq. (5) in the main text with $p=0$, which violates TMD factorisation as shown in the main text.
B.2.2 1.2.2 $k_{t}$-like algorithm from Catani et al. (1992)
In Catani et al. (1992), Catani, Dokshitzer and Webber extended the inclusive $k_{t}$ algorithm initially defined in $e^{+}e^{-}$ annihilation to the DIS case, using the distance measure
$d_{ij}=2\textrm{min}(E_{i}^{2},E_{j}^{2})\frac{1-\cos\theta_{ij}}{R^{2}}\,,\quad d_{iB}=2E_{i}^{2}(1-\cos\theta_{i})$ (20)
computed in the Breit frame  [^7] We recall that the Breit frame is the Lorentz frame in which the virtual photon is a standing wave along the longitudinal ($z$) direction: $q^{\mu}=\delta^{\mu z}Q$ or, in LC notations, $q^{+}=-q^{-}=Q/\sqrt{2}$.. It is clear that this distance measure is not boost invariant along the $\gamma^{*}A$ direction. In order to compute $d_{ij}$ and $d_{iB}$ in the dipole frame, we write these two quantities in terms of Lorentz invariant quantities, using
$E_{i}=\frac{k_{i}\cdot\hat{n}}{Q}$ (21)
with the 4-vector $\hat{n}\equiv 2x_{\rm Bj}P+q$ which reduces to $\hat{n}=(Q,0,0,0)$ in Breit frame and obeys $\hat{n}^{2}=Q^{2}$. Using $M_{ij}^{2}=2E_{i}E_{j}(1-\cos\theta_{ij})$, we have
$d_{ij}=\frac{M_{ij}^{2}}{R^{2}}\textrm{min}\left(\frac{k_{i}\cdot\hat{n}}{k_{j}\cdot\hat{n}},\frac{k_{j}\cdot\hat{n}}{k_{i}\cdot\hat{n}}\right)\,,\quad d_{iB}=\frac{2(k_{i}\cdot\hat{n})(k_{i}\cdot P)}{P\cdot\hat{n}}$ (22)
In the dipole frame, these relations yield
$k_{i}\cdot\hat{n}$ $=\frac{z_{i}Q^{2}}{2}\left[1+\frac{\bm{k}_{i\perp}^{2}}{z_{i}^{2}Q^{2}}\right]\,,\quad d_{iB}=z_{i}^{2}Q^{2}\left[1+\frac{\bm{k}_{i\perp}^{2}}{z_{i}^{2}Q^{2}}\right]$ (23)
Let us now consider the NLO clustering condition $d_{ij}<\textrm{min}(d_{iB},d_{jB})$ in the collinear limit where one can approximate $\bm{k}_{i\perp}\approx({z_{i}}/{z})\bm{P}_{\perp}$. (Like before, $\bm{P}_{\perp}$ and $z$ are the jet transverse momentum and longitudinal momentum fraction in the E-scheme.) A straightforward calculation gives the following clustering condition in the dipole frame:
$\frac{M_{ij}^{2}}{z_{i}z_{j}Q^{2}R^{2}}\left[1+\frac{P_{\perp}^{2}}{z^{2}Q^{2}}\right]^{-1}\leq 1$ (24)
up to corrections suppressed by powers of $R^{2}$. This condition can be further simplified in the regime $Q^{2}\gg P_{\perp}^{2}$ dominated by aligned jet configuration where either $1-z\sim P_{\perp}^{2}/Q^{2}$ or $z\sim P_{\perp}^{2}/Q^{2}$. To leading power in the ratio $P_{\perp}^{2}/Q^{2}$, we can replace Eq. (24) by
$\frac{M_{ij}^{2}}{z_{i}z_{j}Q^{2}R^{2}}\leq\frac{1}{z}\,.$ (25)
B.2.3 1.2.3 $e^{+}e^{-}$ generalised-$k_{t}$ algorithms applied to DIS in the Breit frame
Another common choice of jet definition in the TMD literature is to use the generalised $k_{t}$ algorithms originally designed for $e^{+}e^{-}$ annihilation, which can be extended to DIS in the Breit frame (see e.g. Gutierrez-Reyes et al. (2018, 2019a, 2019b)). Inclusive generalised-$k_{t}$ algorithms in DIS are defined with the spherically invariant distance measure Cacciari et al. (2012)
$d_{ij}=\textrm{min}(E_{i}^{2k},E_{j}^{2k})\frac{1-\cos\theta_{ij}}{1-\cos R}\,,\quad d_{iB}=E_{i}^{2k}$ (26)
where energies and angles are defined in the Breit frame. Note that the distance measures are also not invariant by longitudinal boost. As in the previous case (recall Eq. (21)), the distance measure can be written with Lorentz invariant quantities as
$d_{ij}=\frac{Q^{2}}{(k_{i}\cdot\hat{n})(k_{j}\cdot\hat{n})}\frac{M_{ij}^{2}}{2(1-\cos R)}\textrm{min}\left(\frac{(k_{i}\cdot\hat{n})^{2k}}{Q^{2k}},\frac{(k_{j}\cdot\hat{n})^{2k}}{Q^{2k}}\right)\,,\quad d_{iB}=\frac{(k_{i}\cdot\hat{n})^{2k}}{Q^{2k}}$ (27)
The condition $d_{ij}\leq\textrm{min}(d_{iB},d_{jB})$ for two particles to be clustered within the same jet then reads
$\frac{4\bm{k}_{ij\perp}^{2}}{z_{i}^{2}z_{j}^{2}Q^{2}}\left[1+\frac{\bm{k}_{i\perp}^{2}}{z_{i}^{2}Q^{2}}\right]^{-1}\left[1+\frac{\bm{k}_{j\perp}^{2}}{z_{j}^{2}Q^{2}}\right]^{-1}\leq 2(1-\cos R)\,,$ (28)
in the dipole frame. In the collinear limit, one can again approximate $\bm{k}_{i\perp}\approx\frac{z_{i}}{z}\bm{P}_{\perp}$ giving the condition
$\frac{4M_{ij}^{2}}{z_{i}z_{j}Q^{2}}\left[1+\frac{P_{\perp}^{2}}{z^{2}Q^{2}}\right]^{-2}\leq R^{2}$ (29)
up to power of $R^{2}$ corrections. Let us consider the two regimes which contribute to leading twist (i) $1-z\sim P_{\perp}^{2}/Q^{2}\ll 1$, (ii) $z\sim P_{\perp}^{2}/Q^{2}\ll 1$. In the first case, we can approximate $1+P_{\perp}^{2}/(z^{2}Q^{2})\approx 1\approx z^{-1}$, while in the case (ii), we have $1+P_{\perp}^{2}/(z^{2}Q^{2})\approx Q^{2}/P_{\perp}^{2}\approx z^{-1}$. Thus, to leading power in $P_{\perp}^{2}/Q^{2}$, the clustering condition Eq. (29) can be written
$\frac{4M_{ij}^{2}}{z_{i}z_{j}Q^{2}R^{2}}\leq\frac{1}{z^{2}}\,.$ (30)
This differs from Eq. (25) in the power of $1/z$ in the r.h.s. (together with an irrelevant numerical factor which can be absorbed into a redefinition of $R$). It furthermore differs from the clustering condition A given by Eq. (17) in the replacement $Q^{2}\leftrightarrow P_{\perp}^{2}$.
B.2.4 1.2.4 New jet distance measure given by Eq. (5) in the main text
We recall here that we propose in the main text the distance measure
$d_{ij}=\textrm{min}(k_{i\perp}^{2k},k_{j\perp}^{2k})\frac{M_{ij}^{2}}{z_{i}z_{j}Q^{2}R^{2}}\,,\quad d_{iB}=k_{i\perp}^{2k}$ (31)
which is longitudinally boost invariant (we fix the parameter $p$ to 1 in Eq. (5) of the Letter as we already know from the discussion in the main text that it is the only consistent choice at NLO).
In order to simultaneously account for (i) the $k_{t}$ algorithm by Catani, Dokshitzer and Webber, (ii) the $e^{+}e^{-}$ generalised-$k_{t}$ algorithm used in the DIS Breit frame, and (iii) our new jet distance measure, we shall use the generalised NLO clustering condition
[…]
As demonstrated above, $\beta=1$ corresponds to the DIS $k_{t}$ algorithm defined in Catani et al. (1992), cf. Eq. (25), $\beta=2$ mimics the $e^{+}e^{-}$ generalised-$k_{t}$ algorithms applied to DIS in the Breit frame, cf. Eq. (30), while $\beta=0$ is our new clustering condition, Eq. (31). We shall refer to Eq. (17) as the clustering condition B. We will compute the SIDIS jet cross-section in the limit $Q^{2}\gg P_{\perp}^{2}$ for any value of $\beta$, although the choice $\beta=0$ is the only physical one, as we will argue in the next section.
Appendix C 2 NLO SIDIS cross-section in the limit $Q^{2}\gg P_{\perp}^{2}\gg Q_{s}^{2}$ from the CGC
In this section, we will study the virtual NLO corrections to the cross-section for jet production in SIDIS, with three main objectives in mind: (i) to illustrate the (strong) dependence of the results upon the choice of a jet definition, as summarised in the previous section, (ii) to demonstrate that the only jet definition which is consistent with TMD factorisation at NLO is the condition (5) in the Letter with the choice $p=1$ (equivalently, this is the clustering condition B, Eq. (32) in the present material, with the choice $\beta=0$), and (ii) to establish the result for the virtual corrections shown in Eq. (12) in the Letter, which indeed corresponds to the “physical” jet definition mentioned at point (ii) above. We recall that the compatibility between the jet definition and TMD factorisation is crucial in order to guarantee that the jet structure which is experimentally measured faithfully reflect the partonic picture of the QCD process in perturbation theory, including the natural parton virtualities.
Before we proceed, it is important to stress that, from a diagrammatic point of view, the “virtual corrections” of interest for us here include not only the genuine virtual diagrams — the one-loop corrections to the amplitude —, but also pieces of the would-be real corrections — the NLO amplitudes involving three partons (a quark, an antiquark and a gluon) in the final state — in the kinematical regime where the gluon jet is not well separated from the quark jet (meaning that the quark and the gluon are combined within a same jet by the jet algorithm). This is important for the present purposes since such (would-be) real NLO corrections are the only ones to be sensitive to the actual jet definition. For the one-loop corrections (and to strict NLO accuracy), the struck quark emerges alone in the final state and is unambiguously identified as the jet by all the clustering conditions.
As implicit in the previous discussion, we shall consider the case where the jet measured in SIDIS is generated by the quark. Our calculation is inclusive in the quark longitudinal momentum fraction $z_{1}$: that is, we shall consider both the situation where the quark is fast ($z_{1}\simeq 1$) in the dipole frame (meaning that the quark is the struck fermion in the Breit frame) and the situation where the quark is slow $z_{1}\ll 1$ (so the struck fermion is the antiquark). At NLO, the measured jet can reduce to the quark alone, or it can also include a gluon, that was radiated by either the quark, or the antiquark. Before we turn to the NLO corrections, we first review in the next subsection the leading order (LO) calculation of the SIDIS cross-section in the CGC/dipole picture and its leading twist approximation at $Q^{2}\gg P_{\perp}^{2}$, in which one recovers the TMD factorisation shown in Eq. (1) in the Letter.
C.1 2.1 Brief review of the LO cross-section
The LO SIDIS cross-section in the Colour Glass Condensate effective theory reads (see e.g. Iancu et al. (2021); Caucal et al. (2024))
[…]
×$\times\bar{Q}^{2}K_{1}(\bar{Q}r_{x^{\prime}y})K_{1}(\bar{Q}r_{xy})\frac{\bm{r}_{x^{\prime}y}\cdot\bm{r}_{xy}}{r_{x^{\prime}y}r_{xy}}\,.$ (33)
where $\bar{Q}^{2}=z(1-z)Q^{2}$ and we define $\bm{r}_{xy}=\bm{x}_{\perp}-\bm{y}_{\perp}$ for any two transverse coordinate vectors $\bm{x}_{\perp}$ and $\bm{y}_{\perp}$. In this expression, the quark is measured with transverse momentum $\bm{P}_{\perp}$ and longitudinal momentum fraction $z$. Physically, $\bm{x}_{\perp}$ and $\bm{x}_{\perp}^{\prime}$ represent the transverse coordinates of the quark in the direct amplitude and the complex conjugate amplitude, respectively, and $\bm{b}_{\perp}=(\bm{x}_{\perp}+\bm{x}_{\perp}^{\prime})/2$ is the quark impact parameter. The antiquark is not measured, so its coordinate is the same, and equal to $\bm{y}_{\perp}$, on both sides of the cut. The alternative situation, where the antiquark is measured but the quark is not, gives an identical contribution and is taken into account by an overall factor 2. If one integrates over the $\bm{P}_{\perp}$ and $z$ (with $z\in[0,1]$), one recovers the fully inclusive DIS cross-section at small $x$ and per unit $\bm{b}_{\perp}$ (see e.g. Ducloué et al. (2020)). The CGC colour structure in Eq. (33) is defined as
$\Xi_{\rm LO}(\bm{x}_{\perp},\bm{y}_{\perp},\bm{x}_{\perp}^{\prime})$ $\equiv\left\langle 1-D_{yx^{\prime}}-D_{xy}+D_{xx^{\prime}}\right\rangle_{x}$ (34)
$=\int\mathrm{d}^{2}\bm{q}_{\perp}\left(e^{i\bm{q}_{\perp}\cdot\bm{r}_{xy}}-1\right)\left(e^{-i\bm{q}_{\perp}\cdot\bm{r}_{x^{\prime}y}}-1\right)\mathcal{D}(x,\bm{q}_{\perp}),$ (35)
where the Wilson-line operator $D_{xy}=\frac{1}{N_{c}}\mathrm{Tr}(V({\bm{x}_{\perp}})V^{\dagger}({\bm{y}_{\perp}}))$ represents the $S$-matrix for the elastic scattering of a quark-antiquark colour dipole. The CCG average $\left\langle...\right\rangle_{x}$ should be performed at the scale $x=x_{\rm Bj}$ as explained in the main text of the letter. In the second line of the above expression, we have written the colour correlator in momentum space for later convenience; in particular,
$\mathcal{D}(x,\bm{q}_{\perp})\equiv\int\frac{\mathrm{d}^{2}\bm{r}_{\perp}}{(2\pi)^{2}}e^{-i\bm{q}_{\perp}\cdot\bm{r}_{\perp}}\frac{1}{N_{c}}\left\langle\mathrm{Tr}(V(\bm{b}_{\perp}+\bm{r}_{\perp}/2)V^{\dagger}(\bm{b}_{\perp}-\bm{r}_{\perp}/2))\right\rangle_{x}\,,$ (36)
where we keep the dependence upon the impact parameter $\bm{b}_{\perp}$ implicit. Using this momentum-space representation, one can see that the integrals over $\bm{r}_{xy}$ and $\bm{r}_{x^{\prime}y}$ in Eq. (33) factorise from each other; after also integrating over $z$, one finds
$\left.\frac{\mathrm{d}\sigma_{\rm CGC}^{\gamma_{\rm T}^{\star}+A\to q+X}}{\mathrm{d}^{2}\bm{b}_{\perp}\mathrm{d}^{2}\bm{P}_{\perp}}\right|_{\rm LO}=\frac{\alpha_{\rm em}e_{f}^{2}N_{c}}{\pi^{2}}\int\mathrm{d}^{2}\bm{q}_{\perp}\mathcal{D}(x,\bm{q}_{\perp})\int_{0}^{1/2}\mathrm{d}z\ [z^{2}+(1-z)^{2}]\,\mathcal{H}_{\rm LO}^{i}(\bm{P}_{\perp},\bm{q}_{\perp},z)\mathcal{H}_{\rm LO}^{i*}(\bm{P}_{\perp},\bm{q}_{\perp},z)$ (37)
with
$\mathcal{H}_{\rm LO}^{i}$ $\equiv\int\frac{\mathrm{d}^{2}\bm{r}_{x^{\prime}y}}{2\pi}\ \ e^{-i\bm{P}_{\perp}\cdot\bm{r}_{x^{\prime}y}}\left(e^{i\bm{q}_{\perp}\cdot\bm{r}_{x^{\prime}y}}-1\right)\frac{\bar{Q}K_{1}(\bar{Q}r_{x^{\prime}y})}{r_{x^{\prime}y}}\bm{r}_{x^{\prime}y}^{i}$ (38)
$=\frac{-i(\bm{P}_{\perp}-\bm{q}_{\perp})^{i}}{(\bm{P}_{\perp}-\bm{q}_{\perp})^{2}+\bar{Q}^{2}}-\frac{-i\bm{P}_{\perp}^{i}}{P_{\perp}^{2}+\bar{Q}^{2}}$ (39)
Note that we have integrated $z$ up to $1/2$ and multiplied the result by 2 thanks to the $z\leftrightarrow 1-z$ symmetry of the integrant. The modulus squared of the LO hard factor is readily obtained as
$\mathcal{H}_{\rm LO}^{i}\mathcal{H}_{\rm LO}^{i*}$ $=\frac{P_{\perp}^{2}}{[P_{\perp}^{2}+\bar{Q}^{2}]^{2}}+\frac{(\bm{P}_{\perp}-\bm{q}_{\perp})^{2}}{\left[(\bm{q}_{\perp}-\bm{P}_{\perp})^{2}+\bar{Q}^{2}\right]}-\frac{2\bm{P}_{\perp}\cdot(\bm{P}_{\perp}-\bm{q}_{\perp})}{[(\bm{P}_{\perp}-\bm{q}_{\perp})^{2}+\bar{Q}^{2}][P_{\perp}^{2}+\bar{Q}^{2}]}$ (40)
So far, Eq. (37) is just a rewriting of Eq. (33). At this point we consider its leading-twist (LT) approximation for $Q^{2}\gg P_{\perp}^{2}$. In this limit, the integral over $z$ is dominated by the lower end point $z=0$, and more precisely by values $z\sim P_{\perp}^{2}/Q^{2}$. To the LT accuracy of interest, the various contributions can be estimated as  [^8] A simple way to derive these asymptotic behaviours is to make the change of variable $u=zQ^{2}$ and then to take $Q\to\infty$ in the upper limit of the $u$ integral.
$\int_{0}^{1/2}\mathrm{d}z\frac{1}{[P_{\perp}^{2}+zQ^{2}]^{2}}\simeq\frac{1}{P_{\perp}^{2}Q^{2}}\,,\quad\int_{0}^{1/2}\mathrm{d}z\frac{1}{[P_{\perp}^{2}+zQ^{2}][(\bm{P}_{\perp}-\bm{q}_{\perp})^{2}+zQ^{2}]}$ $\simeq\frac{\ln\left(P_{\perp}^{2}/(\bm{P}_{\perp}-\bm{q}_{\perp})^{2}\right)}{\big{[}P_{\perp}^{2}-(\bm{P}_{\perp}-\bm{q}_{\perp})^{2}\big{]}Q^{2}}\,,$ (41)
thus finally yielding the following, TMD-factorised, result:
$\left.\frac{\mathrm{d}\sigma^{\gamma_{\rm T}^{\star}+A\to q+X}}{\mathrm{d}^{2}\bm{P}_{\perp}}\right|_{\rm LO}$ $=\frac{8\pi^{2}\alpha_{\rm em}e_{f}^{2}}{Q^{2}}\times x\mathcal{F}_{q}(x,\bm{P}_{\perp})\left[1+\mathcal{O}\left(\frac{P_{\perp}^{2}}{Q^{2}}\right)\right]$ (42)
with the (sea) quark TMD in momentum space defined as
$x\mathcal{F}_{q}(x,\bm{P}_{\perp})\equiv\frac{N_{c}}{\pi^{2}}\int\mathrm{d}^{2}\bm{b}_{\perp}\int\frac{\mathrm{d}^{2}\bm{q}_{\perp}}{(2\pi)^{2}}\ \mathcal{D}(x,\bm{q}_{\perp})\left[1-\frac{\bm{P}_{\perp}\cdot(\bm{P}_{\perp}-\bm{q}_{\perp})}{(P_{\perp}^{2}-(\bm{P}_{\perp}-\bm{q}_{\perp})^{2})}\,\ln\frac{P_{\perp}^{2}}{(\bm{P}_{\perp}-\bm{q}_{\perp})^{2}}\right]$ (43)
Eq. (42) agrees with the results from the literature Marquet et al. (2009). As it should be clear from its above derivation, it applies for relatively hard DIS with $Q^{2}\gg P_{\perp}^{2},Q_{s}^{2}$ (the corrections to it are suppressed by powers of $P_{\perp}^{2}/Q^{2}$ and $Q_{s}^{2}/Q^{2}$). However, the saturation “higher-twists”, which describe multiple scattering with $P_{\perp}\sim Q_{s}$ and are proportional to powers of $Q_{s}^{2}/P_{\perp}^{2}$, are resummed to all orders within the TMD. In particular, the typical values of $q_{\perp}$ are of order $Q_{s}$ since for much larger values $q_{\perp}\gg Q_{s}$, $\mathcal{D}(x,\bm{q}_{\perp})$ is rapidly decreasing, as $1/\bm{q}_{\perp}^{4}$ Iancu and Venugopalan (2003). In the ”dilute” limit $P_{\perp}^{2}\gg Q_{s}^{2}$, where one can also neglect the saturation effects, we can simplify the expression of the quark TMD by expanding the quantity in the square bracket in Eq. (43) for $q_{\perp}\sim Q_{s}\ll P_{\perp}$:
$x\mathcal{F}_{q}(x,\bm{P}_{\perp})$ $=\frac{\alpha_{s}}{2\pi^{2}}\frac{1}{P_{\perp}^{2}}\frac{1}{3}\int_{0}^{P_{\perp}^{2}}\mathrm{d}^{2}\bm{q}_{\perp}\ x\mathcal{G}_{D}(x,\bm{q}_{\perp})\,,\quad x\mathcal{G}_{D}(x,\bm{q}_{\perp})=\frac{N_{c}\bm{q}_{\perp}^{2}}{2\pi^{2}\alpha_{s}}\int\mathrm{d}^{2}\bm{b}_{\perp}\mathcal{D}(x,\bm{q}_{\perp})$ (44)
where we have recognised the so-called gluon dipole TMD $x\mathcal{G}_{D}(x,\bm{q}_{\perp})$ Dominguez et al. (2011). The $1/3$ coefficient physically comes from the integral of the $g\to q\bar{q}$ splitting function. Indeed, Eq. (44) can be equivalently rewritten, in the small $x$ regime, in term of the gluon PDF $xG(x,P_{\perp}^{2})$:
$x\mathcal{F}_{q}(x,\bm{P}_{\perp})=\frac{\alpha_{s}}{2\pi^{2}}\frac{1}{P_{\perp}^{2}}\int_{0}^{1}\mathrm{d}\xi\ P_{qg}(\xi)\frac{x}{\xi}G\left(\frac{x}{\xi},P_{\perp}^{2}\right)\,,\quad xG(x,P_{\perp}^{2})=\int_{0}^{P_{\perp}^{2}}\mathrm{d}^{2}\bm{q}_{\perp}\ x\mathcal{G}_{D}(x,\bm{q}_{\perp}).$ (45)
At tree-level, $\mathcal{D}(x,\bm{q}_{\perp})$ has no $x$ dependence, so the integral over $\xi$ can be performed analytically to yield
$\int_{0}^{1}\mathrm{d}\xi\ P_{qg}(\xi)=\int_{0}^{1}\mathrm{d}\xi\ T_{R}\left[\xi^{2}+(1-\xi)^{2}\right]=\frac{1}{3}\,,$ (46)
in agreement with Eq. (44). The emergence of the DGLAP splitting function has a simple interpretation: in order to produce a (sea) $q\bar{q}$ pair with large transverse momentum $P_{\perp},$ a small-$x$ gluon with relatively small transverse momentum $q_{\perp}\ll P_{\perp}$ from the target must undergo a hard splitting.
C.2 2.2 NLO diagrams with poles in transverse dim. reg.
The subsequent calculation of NLO corrections to SIDIS will heavily rely on previously published NLO calculations for dijets (or dihadrons) in DIS Caucal et al. (2021); Taels et al. (2022); Bergabo and Jalilian-Marian (2022); Iancu and Mulian (2023); Bergabo and Jalilian-Marian (2023a) and for jets (or hadrons) in SIDIS Bergabo and Jalilian-Marian (2023b); Caucal et al. (2024), that we shall simply adapt to the kinematics and the jet definitions of relevance for us here. In particular, with respect to Caucal et al. (2024), we re-derive the real contribution to the SIDIS cross-section using the clustering condition B (in Caucal et al. (2024), we only used the clustering condition A) and we extract the leading-twist contribution in the limit $Q^{2}\gg P_{\perp}^{2}$.
Figure 2: Feynman diagrams contributing to the NLO amplitude. The quark-antiquark exchanged diagrams, named with an additional line over the diagram’s label, are not shown. Figure from Caucal et al. (2021).
The Feynman graphs encoding the relevant NLO corrections are illustrated in Fig. 2 at the amplitude level. The two graphs in the first line describe real gluon emissions, while all the other six graphs are virtual corrections — three quark self-energies and three vertex corrections in which the gluon is exchanged between the quark and the antiquark. We only display the real diagrams R1 and R2 where the gluon is emitted by the quark. Clearly, there are corresponding graphs, to be denoted with a bar ($\overline{\rm R1}$, $\overline{\rm R2}$), in which the gluon is emitted by the antiquark. Similarly, there are three antiquark self-energies ($\overline{\rm SE1}$, etc) and an additional vertex correction $\overline{\rm V1}$ where the gluon is emitted by the quark prior to the scattering with the shockwave (SW) and absorbed by the antiquark after the SW. When computing the cross-section, the NLO real graphs are multiplied with each other, while the NLO virtual graphs are multiplied with the LO amplitude. For definiteness, we shall consider the situation where the jet measured by SIDIS has been initiated by the quark — that is, the antiquark and the gluon are integrated over. This in particular implies that the gluon emissions by the antiquark in the final state (after crossing the SW) cannot contribute to the cross-section: their contributions must cancel between real and virtual corrections. In dimensional regularisation, where the final-state self-energies are identically zero (this applies to both SE3 and $\overline{\rm SE3}$), one also has $\overline{\rm R2}\times\overline{\rm R2}^{*}=0$.
When using a jet definition which is infrared and collinear safe (IRC), the cross-section for SIDIS jets computed in the CGC effective theory is free of infrared and collinear divergences, by construction. Yet, such divergences can appear in the individual diagrams and they are only guaranteed to cancel after all the diagrams are combined together. In this section, we display the results for the contributions which have both a $1/\varepsilon$ divergence in transverse dimensional regularisation and logarithmic divergences in the rapidity infrared cut-off $\Lambda^{+}$. We denote $z_{0}\equiv\Lambda^{+}/q^{+}$.
[…]
$\left.\frac{\mathrm{d}\sigma_{\rm CGC}^{\gamma^{*}_{\rm T}+A\to q+X}}{\mathrm{d}^{2}\bm{b}_{\perp}\mathrm{d}^{2}\bm{P}_{\perp}\mathrm{d}z}\right|_{V,\rm pole}$ $=\left.\frac{\mathrm{d}\sigma_{\rm CGC}^{\gamma^{*}_{\rm T}+A\to q+X}}{\mathrm{d}^{2}\bm{b}_{\perp}\mathrm{d}^{2}\bm{P}_{\perp}\mathrm{d}z}\right|_{\rm LO}\times\frac{\alpha_{s}C_{F}}{\pi}\left\{\ln\left(\frac{z}{1-z}\right)\left[\frac{2}{\varepsilon}+\ln\left(e^{\gamma_{E}}\pi\mu^{2}r_{xy}r_{x^{\prime}y}\right)\right]+\frac{1}{2}\ln^{2}\left(\frac{z}{1-z}\right)-\frac{\pi^{2}}{6}+\frac{5}{2}\right.$
$\left.+\left(\ln\left(\frac{1-z}{z_{0}}\right)-\frac{3}{4}\right)\left(\frac{2}{\varepsilon}+\ln\left(e^{\gamma_{E}}\pi\mu^{2}\bm{r}_{xx^{\prime}}^{2}\right)\right)-\frac{1}{4}\right\}$ (47)
with a slight abuse of notation since the quantity inside the curly bracket should appear inside the transverse coordinate integration in the LO cross-section given by Eq. (33).
[…]
As explained at length at the beginning of this section, we are primarily interested in the contributions of type (i), i.e. in the situations where the gluon cannot be distinguished from the quark jet (“in-cone emissions”), which effectively provide virtual corrections to the cross-section for quark jet production. However, it turns out that the separation between in-cone and out-of-cone emissions introduces spurious rapidity divergences, which only cancel after the two types of contributions are added with each other. After this cancellation, we are left with logarithmically-enhanced contributions of the Sudakov type, which are important for us here. To correctly compute these logarithms, we shall include in our calculation the contributions of type (ii) (gluon emissions outside the quark jet) which are needed in order to cancel the spurious rapidity divergences. For that purpose, it is sufficient to consider the situation where the quark jet is measured and the gluon jet is integrated out. The other case, where the gluon jet is measured and the quark one is integrated out, will be not considered.
The contribution to the cross-section from case (i) (a gluon emitted inside the quark jet) can be computed as
$\left.\frac{\mathrm{d}\sigma_{\rm CGC}^{\gamma^{*}_{\rm T}+A\to q+X}}{\mathrm{d}^{2}\bm{b}_{\perp}\mathrm{d}^{2}\bm{P}_{\perp}}\right|_{\rm in}=\int_{0}^{1}\mathrm{d}z$ $\mu^{\varepsilon}\int\mathrm{d}^{2-\varepsilon}\bm{k_{g\perp}}\int\mathrm{d}^{2}\bm{k_{1\perp}}\int_{z_{0}}^{1}\mathrm{d}z_{g}\int_{0}^{1}\mathrm{d}z_{1}\ \delta(z-(z_{1}+z_{g}))\delta(\bm{P}_{\perp}-(\bm{k_{1\perp}}+\bm{k_{g\perp}}))$
×$\times\left.\frac{\mathrm{d}\sigma_{\rm CGC}^{\gamma^{*}_{\rm T}+A\to qg+X}}{\mathrm{d}^{2}\bm{b}_{\perp}\mathrm{d}^{2}\bm{k_{1\perp}}\mathrm{d}^{2}\bm{k_{g\perp}}\mathrm{d}z_{1}\mathrm{d}z_{g}}\right|_{\rm R2\times R2^{*}}\Theta_{\rm in}^{A/B}(\bm{k_{1\perp}},\bm{k_{g\perp}},z_{1},z_{g})$ (48)
where the step function $\Theta_{\rm in}^{A/B}$ enforces the clustering condition $A$ or $B$ respectively given by Eq. (17) and Eq. (32). In terms of the variables $\bm{k_{1\perp}}$, $\bm{k_{g\perp}}$, $z_{1}$ and $z_{g}$, they read
[…]
$\left.\frac{\mathrm{d}\sigma_{\rm CGC}^{\gamma^{*}_{\rm T}+A\to qg+X}}{\mathrm{d}^{2}\bm{b}_{\perp}\mathrm{d}^{2}\bm{k_{1\perp}}\mathrm{d}^{2}\bm{k_{g\perp}}\mathrm{d}z_{1}\mathrm{d}z_{g}}\right|_{\rm R2\times R2^{*}}=\frac{\alpha_{\rm em}e_{f}^{2}N_{c}}{(2\pi)^{6}}\alpha_{s}C_{F}\int\mathrm{d}^{2}\bm{r}_{xy}\mathrm{d}^{2}\bm{r}_{x^{\prime}y}\ \Xi_{\rm LO}(\bm{x}_{\perp},\bm{y}_{\perp},\bm{x}_{\perp}^{\prime})\frac{e^{-i(\bm{k_{g\perp}}+\bm{k_{1\perp}})\cdot\bm{r}_{xx^{\prime}}}}{(\bm{k_{g\perp}}-\frac{z_{g}}{z_{1}}\bm{k_{1\perp}})^{2}}$
×$\times\left\{8\bar{Q}_{\rm R2}^{2}\left[(1-z_{1}-z_{g})^{2}+(z_{1}+z_{g})^{2}\right]\left(1+\frac{z_{g}}{z_{1}}+\frac{z_{g}^{2}}{2z_{1}^{2}}\right)\frac{\bm{r}_{xy}\cdot\bm{r}_{x^{\prime}y}}{r_{xy}r_{x^{\prime}y}}K_{1}(\bar{Q}_{\mathrm{R}2}r_{xy})K_{1}(\bar{Q}_{\mathrm{R}2}r_{x^{\prime}y})\right\}\,,$ (51)
with $\bar{Q}_{\rm R2}^{2}=(1-z_{1}-z_{g})(z_{1}+z_{g})Q^{2}$. This expression exhibits a collinear singularity at $\bm{k_{g\perp}}=({z_{g}}/{z_{1}})\bm{k_{1\perp}}$.
When the above expression is inserted in the integral in Eq. (48), the collinear singularity becomes a pole in dimensional regularisation. Moreover, the presence of the step function $\Theta_{\rm in}^{A/B}$ introduces a double logarithmic divergence in the rapidity cut-off $z_{0}$. The latter comes from the phase-space where both $z_{g}$ and $\bm{k_{g\perp}}$ are small, so it cancels against the contribution associated with the case (ii) where the quark and gluon form two separated jets and the gluon is integrated out Taels et al. (2022). This contribution (gluon emission outside the measured quark jet) reads
$\left.\frac{\mathrm{d}\sigma_{\rm CGC}^{\gamma^{*}_{\rm T}+A\to q+X}}{\mathrm{d}^{2}\bm{b}_{\perp}\mathrm{d}^{2}\bm{P}_{\perp}}\right|_{\rm out}=\int_{0}^{1}\mathrm{d}z$ $\int\mathrm{d}^{2}\bm{k_{g\perp}}\int\mathrm{d}^{2}\bm{k_{1\perp}}\int_{z_{0}}^{1}\mathrm{d}z_{g}\int_{0}^{1}\mathrm{d}z_{1}\ \delta(z-z_{1})\delta(\bm{P}_{\perp}-\bm{k_{1\perp}})$
×$\times\left.\frac{\mathrm{d}\sigma_{\rm CGC}^{\gamma^{*}_{\rm T}+A\to qg+X}}{\mathrm{d}^{2}\bm{b}_{\perp}\mathrm{d}^{2}\bm{k_{1\perp}}\mathrm{d}^{2}\bm{k_{g\perp}}\mathrm{d}z_{1}\mathrm{d}z_{g}}\right|_{\rm R2\times R2^{*}}\left(1-\Theta_{\rm in}^{A/B}(\bm{k_{1\perp}},\bm{k_{g\perp}},z_{1},z_{g})\right)\,.$ (52)
After combining the real corrections — the ”in” piece in Eq. (48) and the ”out” piece in Eq. (52) — with the virtual corrections in Eq. (47) and by following the same steps as in e.g. Caucal et al. (2021); Taels et al. (2022); Caucal et al. (2022) we find (with the same slight abuse of notation as in Eq. (47))
$\left.\frac{\mathrm{d}\sigma_{\rm CGC}^{\gamma^{*}_{\rm T}+A\to q+X}}{\mathrm{d}^{2}\bm{b}_{\perp}\mathrm{d}^{2}\bm{P}_{\perp}}\right|_{B\rm-def}$ $=\int_{0}^{1}\mathrm{d}z\left.\frac{\mathrm{d}\sigma_{\rm CGC}^{\gamma^{*}_{\rm T}+A\to q+X}}{\mathrm{d}^{2}\bm{b}_{\perp}\mathrm{d}^{2}\bm{P}_{\perp}\mathrm{d}z}\right|_{\rm LO}\times\frac{\alpha_{s}C_{F}}{\pi}\left\{-\frac{1}{2}\ln^{2}(1-z)+\frac{3-2\beta}{2}\ln^{2}(z)-\frac{3(1-\beta/2)}{2}\ln(z)\right.$
$-\frac{3}{4}\ln\left(\frac{Q^{2}r_{xx^{\prime}}^{2}R^{2}}{c_{0}^{2}}\right)+\frac{11}{2}-\frac{\pi^{2}}{2}+\ln\left(\frac{z}{1-z}\right)\ln\left(\frac{Q^{2}r_{xy}r_{x^{\prime}y}R^{2}}{c_{0}^{2}}\right)-(1-\beta)\ln(z)\ln(1-z)$
$\left.+\int_{0}^{1-z}\frac{\mathrm{d}z_{g}}{z_{g}}\ln\left(\frac{Q^{2}r_{xx^{\prime}}^{2}R^{2}z_{g}^{2}}{c_{0}^{2}z^{\beta}}\right)\left[1-J_{0}\left(\frac{z_{g}P_{\perp}r_{xx^{\prime}}}{z}\right)\right]+\mathcal{O}(z)+\mathcal{O}(1-z)+\mathcal{O}(R^{2})\right\}$ (53)
for the clustering condition $B$ which depends on the parameter $\beta$ (see Eq. (32)). As usual in this context, we include the constant $c_{0}=2e^{-\gamma_{E}}$ with $\gamma_{E}$ the Euler-Mascheroni number in the arguments of the logarithms. In this expression, we have discarded powers of $z$ and $1-z$ inside the $z$ integration as they give contributions suppressed by powers of $P_{\perp}^{2}/Q^{2}$ after performing the integral over $z$. On the other hand, for the clustering condition $A$, we have
$\left.\frac{\mathrm{d}\sigma_{\rm CGC}^{\gamma^{*}_{\rm T}+A\to q+X}}{\mathrm{d}^{2}\bm{b}_{\perp}\mathrm{d}^{2}\bm{P}_{\perp}}\right|_{A\rm-def}=\int_{0}^{1}\mathrm{d}z\left.\frac{\mathrm{d}\sigma_{\rm CGC}^{\gamma^{*}_{\rm T}+A\to q+X}}{\mathrm{d}^{2}\bm{b}_{\perp}\mathrm{d}^{2}\bm{P}_{\perp}\mathrm{d}z}\right|_{\rm LO}\times\frac{\alpha_{s}C_{F}}{\pi}\left\{-\frac{1}{2}\ln^{2}\left(\frac{z}{1-z}\right)+\ln\left(\frac{z}{1-z}\right)\ln\left(\frac{P_{\perp}^{2}r_{xy}r_{x^{\prime}y}R^{2}}{c_{0}^{2}}\right)+7\right.$
$\left.-\frac{2\pi^{2}}{3}-\frac{3}{4}\ln\left(\frac{4P_{\perp}^{2}r_{xx^{\prime}}^{2}R^{2}}{c_{0}^{2}}\right)+\int_{0}^{1-z}\frac{\mathrm{d}z_{g}}{z_{g}}\ln\left(\frac{P_{\perp}^{2}r_{xx^{\prime}}^{2}R^{2}z_{g}^{2}}{c_{0}^{2}z^{2}}\right)\left[1-J_{0}\left(\frac{z_{g}P_{\perp}r_{xx^{\prime}}}{z}\right)\right]+\mathcal{O}(z)+\mathcal{O}(1-z)+\mathcal{O}(R^{2})\right\}$ (54)
Importantly, for both clustering conditions ($A$ or $B$), all soft and collinear divergences have cancelled in the sum of real and virtual corrections, as they should (since we employ IRC safe jet definitions): indeed, there is no $1/\varepsilon$ pole in Eqs. (53)–(54), nor $\ln^{2}z_{0}$, or $\ln z_{0}$, divergences. However, the two clustering conditions give different Sudakov double and single logarithms, as we shall shortly see.
So far, our calculation was quite general, in the sense that Eqs. (53)–(54) holds for generic values of the transverse momentum $P_{\perp}$ of the measured jet. Remember however that (a) TMD factorisation is expected only to leading order in the leading twist approximation at $P_{\perp}^{2}\ll Q^{2}$ and (b) to extract the (double and single) Sudakov logarithms it suffices to consider the dilute limit at $P_{\perp}^{2}\gg Q_{s}^{2}$. So, in what follows we shall simplify the above results via approximations which are appropriate in the range $Q^{2}\gg P_{\perp}^{2}\gg Q_{s}^{2}$.
To extract the leading twist contributions of Eqs. (53)–(54), i.e. their dominant contributions in the expansion in powers of $P_{\perp}^{2}/Q^{2}$, we follow the same steps as in the calculation of the LO cross-section in the previous subsection. One should note that some terms inside the curly bracket contribute either via the endpoint $z=1$, or via $z=0$ (more precisely, $z\sim P_{\perp}^{2}/Q^{2}\ll 1$), or through both. For example, the term expressed as an integral over $z_{g}$ only contributes via the endpoint $z=0$, since it vanishes as $z\to 1$. In the limit $z\to 0$, one can further approximate its contribution as
$\int_{0}^{1-z}\frac{\mathrm{d}z_{g}}{z_{g}}\ln\left(\frac{Q^{2}r_{xx^{\prime}}^{2}R^{2}z_{g}^{2}}{c_{0}^{2}z^{\beta}}\right)\left[1-J_{0}\left(\frac{z_{g}P_{\perp}r_{xx^{\prime}}}{z}\right)\right]$ $\simeq\int_{\frac{c_{0}z}{P_{\perp}r_{xx^{\prime}}}}^{1}\frac{\mathrm{d}z_{g}}{z_{g}}\ln\left(\frac{Q^{2}r_{xx^{\prime}}^{2}R^{2}z_{g}^{2}}{c_{0}^{2}z^{\beta}}\right)$
$=\frac{1}{2}\ln\left(\frac{P_{\perp}^{2}r_{xx^{\prime}}^{2}}{z^{2}c_{0}^{2}}\right)\left[\ln\left(\frac{Q^{2}r_{xx^{\prime}}^{2}R^{2}z^{1-\beta}}{c_{0}^{2}}\right)-\frac{1}{2}\ln\left(\frac{P_{\perp}^{2}r_{xx^{\prime}}^{2}}{c_{0}^{2}}\right)\right]$ (55)
up to corrections suppressed by powers of $z\sim P_{\perp}^{2}/Q^{2}$. Recall that $z$ is the longitudinal fraction of the measured jet, that is, either the quark alone ($z=z_{1}$), or the quark accompanied by the gluon ($z=z_{1}+z_{g}$). So, the fact that $z\ll 1$ means that the fermion struck by the photon is the unmeasured antiquark ($z_{2}\simeq 1$). This in turn implies that, in the Breit frame, the measured jet propagates in the fragmentation region of the nuclear target. This is interesting since, as we shall shortly see, the contribution in Eq. (55) leads to a Sudakov double logarithm. The integral over $z_{g}$ in Eq. (54) can be similarly estimated: its result can be inferred from Eq. (55) by replacing $Q^{2}\to P_{\perp}^{2}$ and $\beta\to 2$.
Furthermore, in the dilute limit $P_{\perp}\gg Q_{s}$, one can approximate $r_{xx^{\prime}}\sim r_{xy}\sim r_{x^{\prime}y}\sim c_{0}/P_{\perp}$ inside the logarithms in Eqs. (53)–(54). To see this, recall the factorised structure of the LO cross-section, cf. Eq. (37): when $P_{\perp}\gg q_{\perp}\sim Q_{s}$, the values of $r_{xy}$ and $r_{x^{\prime}y}$ are fixed by Fourier transforms like Eq. (39) to values of order $1/P_{\perp}$; the same is true for $r_{xx^{\prime}}=|\bm{r}_{xy}-\bm{r}_{x^{\prime}y}|$, cf. Eq. (33). After also integrating over $z$, we find
$\left.\frac{\mathrm{d}\sigma^{\gamma^{*}_{\rm T}+A\to q+X}}{\mathrm{d}^{2}\bm{b}_{\perp}\mathrm{d}^{2}\bm{P}_{\perp}}\right|_{A\rm-def}$ $=\left.\frac{\mathrm{d}\sigma^{\gamma^{*}_{\rm T}+A\to q+X}}{\mathrm{d}^{2}\bm{b}_{\perp}\mathrm{d}^{2}\bm{P}_{\perp}}\right|_{\rm LO}\times\frac{\alpha_{s}C_{F}}{\pi}\left[-\ln(R)\ln\left(\frac{Q^{2}}{P_{\perp}^{2}}\right)-\frac{3}{2}\ln(2R)+7-\frac{2\pi^{2}}{3}\right]$ (56)
$\left.\frac{\mathrm{d}\sigma^{\gamma^{*}_{\rm T}+A\to q+X}}{\mathrm{d}^{2}\bm{b}_{\perp}\mathrm{d}^{2}\bm{P}_{\perp}}\right|_{B\rm-def}$ $=\left.\frac{\mathrm{d}\sigma^{\gamma^{*}_{\rm T}+A\to q+X}}{\mathrm{d}^{2}\bm{b}_{\perp}\mathrm{d}^{2}\bm{P}_{\perp}}\right|_{\rm LO}\times\frac{\alpha_{s}C_{F}}{\pi}\left[\frac{1}{2}\ln^{2}\left(\frac{Q^{2}}{P_{\perp}^{2}}\right)+\left(\ln(R)-\frac{3\beta}{8}\right)\ln\left(\frac{Q^{2}}{P_{\perp}^{2}}\right)-\frac{3}{2}\ln(R)+\frac{11}{2}-\frac{\pi^{2}}{2}\right]$ (57)
up to corrections suppressed by powers of $P_{\perp}^{2}/Q^{2}$ and $Q_{s}^{2}/P_{\perp}^{2}$.
By inspection of the above results, one can immediately notice the crucial role of the jet definition for the value of the Sudakov logarithms in the NLO impact factor. For the clustering condition $A$, the sum of the diagrams with a $1/\varepsilon$ divergence does not generate any Sudakov double logarithm, in sharp contrast with the clustering condition $B$. Physically this difference can be understood as follows. The double logarithmic contribution in Eq. (57) comes from real gluon emissions which remain inside the quark jet  [^9] Indeed, one can easily check that the double logarithmic piece in the r.h.s. of Eq. (57) has been generated by the first term in the r.h.s. Eq. (55) (the second term there vanishes after identifying $r_{xx^{\prime}}\simeq c_{0}/P_{\perp}$)., cf. Eq. (48), and which are relatively hard: the transverse momenta of the quark and the gluon obey $Q^{2}\gg k_{1\perp}^{2}\simeq k_{g\perp}^{2}\gg P_{\perp}^{2}$ (with $\bm{k_{1\perp}}+\bm{k_{g\perp}}=\bm{P}_{\perp}$, of course). Such hard in-cone emissions are allowed by the clustering condition $B$, but not also by condition $A$. To see this, notice that in the hard kinematics at hand, the quark-gluon relative momentum is itself hard: $(z_{g}\bm{k_{1\perp}}-z_{1}\bm{k_{g\perp}})^{2}\sim k_{1\perp}^{2}\sim k_{g\perp}^{2}$. Accordingly, it is only the second $\theta$–function in Eqs. (49)–(50), namely $\Theta_{\rm in}^{B}$, which has a non-zero support in this kinematic range. The difference between the two above results is indeed crucial for what follows. All the other NLO corrections that we shall compute in what follows turn out to be independent of the clustering condition, so the dissymmetry between Eq. (56) and, respectively, Eq. (57) will transmit to the final results. The above argument also suggests that the double logarithmic correction visible in Eq. (57) should have no physical consequences: this is associated with in-cone gluon emissions, which do not modify the final state. And indeed, we shall later find that this contribution is compensated by virtual emissions with similar kinematics. But the effect of the latter is independent of the clustering condition, as just mentioned. Since the (physically expected) real versus virtual compensation occurs for the clustering condition $B$ alone, this discussion already suggests that the clustering condition $A$ leads to unphysical features. We shall return to this point in our summary in Sect. 2.5.
Even for the case of the clustering condition $B$, we observe a dependence upon the parameter $\beta$ entering at single logarithmic accuracy. After combining Eq. (57) with the remaining diagrams (those which are finite when $\varepsilon\to 0$), we will see that only the choice $\beta=0$ in the clustering condition $B$ gives a result which is consistent with TMD factorisation.
C.3 2.3 Virtual diagrams without poles
[…]
C.3.1 2.3.1 Dressed self-energy correction on the quark
[…]
$\left.\frac{\mathrm{d}\sigma_{\rm CGC}^{\gamma_{\rm T}^{\star}+A\to q+X}}{\mathrm{d}^{2}\bm{b}_{\perp}\mathrm{d}^{2}\bm{P}_{\perp}\mathrm{d}z}\right|_{\rm SE1\times LO^{*}}$ $=\frac{\alpha_{\rm em}e_{f}^{2}N_{c}}{(2\pi)^{4}}\int\mathrm{d}^{2}\bm{r}_{xy}\mathrm{d}^{2}\bm{r}_{x^{\prime}y}\frac{\mathrm{d}^{2}\bm{r}_{zx}}{\bm{r}_{zx}^{2}}e^{-i\bm{P}_{\perp}\cdot\bm{r}_{xx^{\prime}}}\frac{\alpha_{s}}{\pi^{2}}2z(z^{2}+(1-z)^{2})\frac{\bar{Q}K_{1}(\bar{Q}r_{x^{\prime}y})\bm{r}_{x^{\prime}y}^{i}}{r_{x^{\prime}y}}$
×$\times\int_{0}^{z}\frac{\mathrm{d}z_{g}}{z_{g}}\Bigg{\{}\left(1-\frac{z_{g}}{z}+\frac{z_{g}^{2}}{2z^{2}}\right)\left[\bar{Q}K_{1}(QX_{V})\frac{\sqrt{1-z}\bm{R}_{\rm SE}^{i}}{X_{V}}e^{-i\frac{z_{g}}{z}\bm{P}_{\perp}\cdot\bm{r}_{zx}}\Xi_{\rm NLO,1}(\bm{x}_{\perp},\bm{y}_{\perp},\bm{z}_{\perp},\bm{x}_{\perp}^{\prime})\right.$
$\left.-\bar{Q}K_{1}(\bar{Q}r_{xy})\exp\left(-\frac{\bm{r}_{zx}^{2}}{\bm{r}_{xy}^{2}e^{\gamma_{E}}}\right)\frac{\bm{r}_{xy}^{i}}{r_{xy}}C_{F}\Xi_{\rm LO}(\bm{x}_{\perp},\bm{y}_{\perp},\bm{x}_{\perp}^{\prime})\right]\Bigg{\}}_{z_{\star}}\,,$ (58)
where we recall that $\bar{Q}^{2}=z(1-z)Q^{2}$ and the additional notations read as follows
$X_{V}^{2}$ $=(1-z)(z-z_{g})\bm{r}_{xy}^{2}+z_{g}(z-z_{g})\bm{r}_{zx}^{2}+(1-z)z_{g}\bm{r}_{zy}^{2}\,,$ (59)
$\bm{R}_{\rm SE}$ $=\bm{r}_{xy}+\frac{z_{g}}{z}\bm{r}_{zx},$ (60)
$\Xi_{\rm NLO,1}(\bm{x}_{\perp},\bm{y}_{\perp},\bm{z}_{\perp},\bm{x}_{\perp}^{\prime})$ $=\frac{N_{c}}{2}\left\langle 1-D_{yx^{\prime}}-D_{xz}D_{zy}+D_{zx^{\prime}}D_{xz}\right\rangle_{x}-\frac{1}{2N_{c}}\left\langle 1-D_{xy}-D_{yx^{\prime}}+D_{xx^{\prime}}\right\rangle_{x}.$ (61)
Notice the subscript $z_{\star}$ on the accolades in Eq. (58): {}$\{\cdots\}_{z_{\star}}$. This notation expresses the fact that the rapidity divergence as $z_{g}\to 0$ has already been subtracted off, using an arbitrary ”rapidity factorisation scale” $z_{\star}=k_{\star}^{+}/q^{+}$. More precisely, our notation means
$\int_{0}^{a}\frac{\mathrm{d}z_{g}}{z_{g}}\left\{f(z_{g})\right\}_{z_{\star}}\equiv\int_{0}^{a}\frac{\mathrm{d}z_{g}}{z_{g}}\left[f(z_{g})-\Theta(z_{\star}-z_{g})f(0)\right]\,,$ (62)
for any function $f(z_{g})$ which has a well-defined limit as $z_{g}\to 0$. As discussed in the main text of the letter, we use $z_{\star}=x_{\star}P_{\perp}^{2}/Q^{2}$ in order to separate the NLO impact factor from the (kinematically constrained) BK evolution of the dipole cross-section. The dependence upon the arbitrary $x_{\star}$ number can be used to gauge the sensitivity of the final result to the arbitrary separation between high energy and DGLAP+Sudakov evolution.
Here comes an important observation: the leading twist piece of the expression Eq. (58) can be obtained from the transverse coordinate domain where $r_{zx}\sim 1/Q\ll r_{xy}\sim 1/P_{\perp}$. Physically, this means that the virtual quark-gluon fluctuation is much harder than the measured quark. Similar to the LO calculation, we shall systematically consider the aligned jet configuration such that $1-z\sim P_{\perp}^{2}/Q^{2}\ll 1$ (the other endpoint $z\to 0$ is power suppressed). In this limit, the colour structure simply reduces to the LO one times a Casimir factor $C_{F}$, namely
$\Xi_{\rm NLO,1}(\bm{x}_{\perp},\bm{y}_{\perp},\bm{z}_{\perp},\bm{x}_{\perp}^{\prime})$ $=C_{F}\left\langle 1-D_{yx^{\prime}}-D_{xy}+D_{xx^{\prime}}\right\rangle_{x}+\mathcal{O}(Q_{s}^{2}\bm{r}_{zx}^{2})$ (63)
$\simeq C_{F}\int\mathrm{d}^{2}\bm{q}_{\perp}\left(e^{i\bm{q}_{\perp}\cdot\bm{r}_{xy}}-1\right)\left(e^{-i\bm{q}_{\perp}\cdot\bm{r}_{x^{\prime}y}}-1\right)\mathcal{D}(x,\bm{q}_{\perp})$ (64)
In Eq. (58), one can also remove the phase $e^{-i\frac{z_{g}}{z}\bm{P}_{\perp}\cdot\bm{r}_{zx}}$ since the argument is of order $P_{\perp}/Q\ll 1$. Finally, the expression of $X_{V}$ in Eq. (59) simplifies as well: in the limit we are considering, $r_{zx}\ll r_{xy}\simeq r_{zy}$, which gives
$X_{V}^{2}\simeq z(1-z)\bm{r}_{xy}^{2}+z_{g}(1-z_{g})\bm{r}_{zx}^{2}\,,\qquad\bm{R}_{\rm SE}\simeq\bm{r}_{xy}.$ (65)
After these simplifications, and integrating over the rapidity $z$ of the jet, Eq. (58) reads Caucal et al. (2021, 2024)
$\left.\frac{\mathrm{d}\sigma^{\gamma_{\rm T}^{\star}+A\to q+X}}{\mathrm{d}^{2}\bm{b}_{\perp}\mathrm{d}^{2}\bm{P}_{\perp}}\right|_{\rm SE1}=\frac{\alpha_{\rm em}e_{f}^{2}N_{c}}{2\pi}\frac{4\alpha_{s}C_{F}}{\pi^{2}}\int\mathrm{d}^{2}\bm{q}_{\perp}\mathcal{D}(x,\bm{q}_{\perp})\int_{0}^{1}\mathrm{d}z\int_{x_{\star}P_{\perp}^{2}/Q^{2}}^{1}\frac{\mathrm{d}z_{g}}{z_{g}}\left(1-z_{g}+\frac{z_{g}^{2}}{2}\right)\mathcal{H}_{\rm LO}^{i*}\mathcal{H}_{\rm SE1}^{i}.$ (66)
[…]
$\mathcal{H}_{\rm SE1}^{i}$ $\equiv\int\frac{\mathrm{d}^{2}\bm{r}_{zx}}{(2\pi)}\int\frac{\mathrm{d}^{2}\bm{r}_{xy}}{(2\pi)}e^{-i\bm{P}_{\perp}\cdot\bm{r}_{xy}}\left(e^{i\bm{q}_{\perp}\cdot\bm{r}_{xy}}-1\right)\frac{\bm{r}_{xy}^{i}}{\bm{r}_{zx}^{2}}\left[\frac{\bar{Q}K_{1}\left(\bar{Q}\sqrt{\bm{r}_{xy}^{2}+\omega\bm{r}_{zx}^{2}}\right)}{\sqrt{\bm{r}_{xy}^{2}+\omega\bm{r}_{zx}^{2}}}-\frac{\bar{Q}K_{1}(\bar{Q}r_{xy})\exp\left(-\frac{\bm{r}_{zx}^{2}}{\bm{r}_{xy}^{2}e^{\gamma_{E}}}\right)}{r_{xy}}\right]$ (67)
with the new notation
$\omega\equiv\frac{z_{g}(1-z_{g})}{z(1-z)}\approx\frac{z_{g}(1-z_{g})}{1-z}$ (68)
[…]
$\mathcal{H}_{\rm SE1}^{i}$ $=\frac{-i(\bm{P}_{\perp}^{i}-\bm{q}_{\perp}^{i})}{2((\bm{P}_{\perp}-\bm{q}_{\perp})^{2}+\bar{Q}^{2})}\left[\ln\left(\frac{1}{\omega}\right)-\frac{\bar{Q}^{2}}{(\bm{P}_{\perp}-\bm{q}_{\perp})^{2}}\ln\left(1+\frac{(\bm{P}_{\perp}-\bm{q}_{\perp})^{2}}{\bar{Q}^{2}}\right)\right]$
$+\frac{i\bm{P}_{\perp}^{i}}{2(P_{\perp}^{2}+\bar{Q}^{2})}\left[\ln\left(\frac{1}{\omega}\right)-\frac{\bar{Q}^{2}}{P_{\perp}^{2}}\ln\left(1+\frac{P_{\perp}^{2}}{\bar{Q}^{2}}\right)\right]$ (69)
Note that this hard factor is leading power, as one easily sees this by comparing with the LO one in Eq. (39) (recall that $\bar{Q}^{2}\sim P_{\perp}^{2}$ for the relevant value $1-z\sim P_{\perp}^{2}/Q^{2}$). The last steps consist in doing the contraction over the index $i$, expanding the result for $P_{\perp}\gg q_{\perp}$ and finally performing the integrals over $z_{g}$ and $z$. The integral over $z$ is performed along the lines of Eq. (41) and gives back the LO cross-section in the limit $P_{\perp}^{2}\gg Q_{s}^{2}$. We thus get
$\left.\frac{\mathrm{d}\sigma^{\gamma_{\rm T}^{\star}+A\to q+X}}{\mathrm{d}^{2}\bm{P}_{\perp}}\right|_{\rm SE1}=\left.\frac{\mathrm{d}\sigma^{\gamma_{\rm T}^{\star}+A\to q+X}}{\mathrm{d}^{2}\bm{P}_{\perp}}\right|_{\rm LO}\times\frac{\alpha_{s}C_{F}}{\pi}$ $\left[-\frac{3}{2}+\frac{\pi^{2}}{6}+\frac{3}{4}\ln\left(\frac{Q^{2}}{P_{\perp}^{2}}\right)-\ln\left(\frac{Q^{2}}{x_{\star}P_{\perp}^{2}}\right)\ln\left(\frac{Q^{2}}{P_{\perp}^{2}}\right)+\frac{1}{2}\ln^{2}\left(\frac{Q^{2}}{x_{\star}P_{\perp}^{2}}\right)\right.$
$\left.-\frac{2}{3}\left(-\frac{3}{4}+\ln\left(\frac{Q^{2}}{x_{\star}P_{\perp}^{2}}\right)\right)\right]$ (70)
To summarise, the (UV-regularised) quark self-energy correction where the gluon crosses the shock-wave — the amplitude SE1 in Fig. 2 — contributes to the Sudakov double logarithm with a coefficient $-\frac{\alpha_{s}C_{F}}{2\pi}\ln^{2}(Q^{2}/P_{\perp}^{2})$. It is also interesting to notice that the double-logarithmic terms of the type
$\ln\left(\frac{Q^{2}}{P_{\perp}^{2}}\right)\,\ln\left(\frac{1}{x_{*}}\right),$ (71)
which would mix the Sudakov logarithms with the logarithm of the arbitrary factorisation scale $x_{*}$, exactly cancel in the r.h.s. of Eq. (70). We shall comment on the meaning of this cancellation in the summary section 2.5.
C.3.2 2.3.2 Dressed vertex correction from the antiquark to the quark
We then consider the vertex correction denoted as V1 in in Fig. 2: the gluon is first emitted by the antiquark, it crosses the shockwave background field and then it is absorbed by the quark. Its contribution to the SIDIS cross-section is Caucal et al. (2024)
$\left.\frac{\mathrm{d}\sigma_{\rm CGC}^{\gamma_{\rm T}^{\star}+A\to q+X}}{\mathrm{d}^{2}\bm{b}_{\perp}\mathrm{d}^{2}\bm{P}_{\perp}\mathrm{d}z}\right|_{\rm V1\times LO^{*}}=\frac{\alpha_{\rm em}e_{f}^{2}N_{c}}{(2\pi)^{4}}\int\mathrm{d}^{2}\bm{r}_{xy}\mathrm{d}^{2}\bm{r}_{x^{\prime}y}\mathrm{d}^{2}\bm{r}_{zx}e^{-i\bm{P}_{\perp}\cdot\bm{r}_{xx^{\prime}}}\frac{\alpha_{s}}{\pi^{2}}2z(1-z)Q\bar{Q}\frac{K_{1}(\bar{Q}r_{x^{\prime}y})}{r_{x^{\prime}y}}\Xi_{\rm NLO,1}(\bm{x}_{\perp},\bm{y}_{\perp},\bm{z}_{\perp},\bm{x}_{\perp}^{\prime})$
×$\times\int_{0}^{z}\frac{\mathrm{d}z_{g}}{z_{g}}\left\{e^{-i\frac{z_{g}}{z}\bm{k}_{\perp}\cdot\bm{r}_{zx}}\frac{K_{1}(QX_{V})}{X_{V}}\left[-\frac{z_{g}(z-z_{g})(1+z_{g}-2z)^{2}}{2z^{2}(1-z)}\frac{(\bm{R}_{\rm V}\times\bm{r}_{x^{\prime}y})(\bm{r}_{zx}\times\bm{r}_{zy})}{\bm{r}_{zx}^{2}\bm{r}_{zy}^{2}}\right.\right.$
$\left.\left.-[z(z-z_{g})+(1-z)(1-z+z_{g})]\left(1-\frac{z_{g}}{z}\right)\left(1+\frac{z_{g}}{1-z}\right)\left(1-\frac{z_{g}}{2z}-\frac{z_{g}}{2(1-z+z_{g})}\right)\frac{(\bm{R}_{\rm V}\cdot\bm{r}_{x^{\prime}y})(\bm{r}_{zx}\cdot\bm{r}_{zy})}{\bm{r}_{zx}^{2}\bm{r}_{zy}^{2}}\right]\right\}_{z_{\star}}$ (72)
where the power-suppressed instantaneous contribution has already been discarded. We employ the same approximations as for the previous diagrams. Namely, for $z$ close to 1, $\bm{R}_{\rm V}=\bm{r}_{xy}-\frac{z_{g}}{1-z+z_{g}}\bm{r}_{zy}\simeq-\bm{r}_{zx}$. In the same limit $1-z\sim P_{\perp}^{2}/Q^{2}\ll 1$, the prefactors depending upon $z$ and $z_{g}$ in front of both terms inside the square bracket become equal to each other, namely they both reduce to
$\frac{-z_{g}(1-z_{g})^{3}}{2(1-z)}$ (73)
so that we can combine together the two vectorial structures
$\frac{(\bm{R}_{\rm V}\times\bm{r}_{x^{\prime}y})(\bm{r}_{zx}\times\bm{r}_{zy})+(\bm{R}_{\rm V}\cdot\bm{r}_{x^{\prime}y})(\bm{r}_{zx}\cdot\bm{r}_{zy})}{\bm{r}_{zx}^{2}\bm{r}_{zy}^{2}}$ $\approx-\frac{(\bm{r}_{zx}\times\bm{r}_{x^{\prime}y})(\bm{r}_{zx}\times\bm{r}_{xy})+(\bm{r}_{zx}\cdot\bm{r}_{x^{\prime}y})(\bm{r}_{zx}\cdot\bm{r}_{xy})}{\bm{r}_{zx}^{2}\bm{r}_{xy}^{2}}$ (74)
$=-\frac{\bm{r}_{xy}\cdot\bm{r}_{x^{\prime}y}}{\bm{r}_{xy}^{2}}$ (75)
In the end, we can cast this diagram (plus its complex conjugate) into the simpler form
$\left.\frac{\mathrm{d}\sigma^{\gamma_{\rm T}^{\star}+A\to q+X}}{\mathrm{d}^{2}\bm{b}_{\perp}\mathrm{d}^{2}\bm{P}_{\perp}}\right|_{\rm V1}=\frac{\alpha_{\rm em}e_{f}^{2}N_{c}}{2\pi}\frac{2\alpha_{s}C_{F}}{\pi^{2}}\int\mathrm{d}^{2}\bm{q}_{\perp}\mathcal{D}(x,\bm{q}_{\perp})\int_{0}^{1}\mathrm{d}z\int_{x_{\star}P_{\perp}^{2}/Q^{2}}^{1}\frac{\mathrm{d}z_{g}}{z_{g}}\frac{z_{g}(1-z_{g})^{3}}{1-z}\mathcal{H}_{\rm LO}^{i*}\mathcal{H}_{\rm V1}^{i}$ (76)
with the hard factors
$\mathcal{H}_{\rm V1}^{i}$ $=\int\frac{\mathrm{d}^{2}\bm{r}_{zx}}{(2\pi)}\int\frac{\mathrm{d}^{2}\bm{r}_{xy}}{(2\pi)}e^{-i\bm{k}_{\perp}\cdot\bm{r}_{xy}}\left(e^{i\bm{q}_{\perp}\cdot\bm{r}_{xy}}-1\right)\frac{\bm{r}_{xy}^{i}}{\bm{r}_{xy}^{2}}\frac{\bar{Q}K_{1}\left(\bar{Q}\sqrt{\bm{r}_{xy}^{2}+\omega\bm{r}_{zx}^{2}}\right)}{\sqrt{\bm{r}_{xy}^{2}+\omega\bm{r}_{zx}^{2}}}$ (77)
$=\frac{(-i)}{2\omega}\left[\frac{\bm{P}_{\perp}^{i}-\bm{q}_{\perp}^{i}}{(\bm{P}_{\perp}-\bm{q}_{\perp})^{2}}\ln\left(1+\frac{(\bm{P}_{\perp}-\bm{q}_{\perp})^{2}}{\bar{Q}^{2}}\right)-\frac{\bm{P}_{\perp}^{i}}{P_{\perp}^{2}}\ln\left(1+\frac{P_{\perp}^{2}}{\bar{Q}^{2}}\right)\right]$ (78)
The overall factor $1/\omega\propto 1-z$ in the r.h.s. of the above expression for $\mathcal{H}_{\rm V1}^{i}$ compensates the would-be divergence at $z\to 1$ of the rational function in Eq. (76). We finally perform the integrals over $z$ and $z_{g}$ and take both the high-$Q^{2}$ limit $Q^{2}\gg P_{\perp}^{2}$ and the dilute limit $P_{\perp}\gg Q_{s}\sim q_{\perp}$ in order to simplify the calculation. This gives:
$\left.\frac{\mathrm{d}\sigma^{\gamma_{\rm T}^{\star}+A\to q+X}}{\mathrm{d}^{2}\bm{P}_{\perp}}\right|_{\rm V1}=\left.\frac{\mathrm{d}\sigma^{\gamma_{\rm T}^{\star}+A\to q+X}}{\mathrm{d}^{2}\bm{P}_{\perp}}\right|_{\rm LO}\times\frac{\alpha_{s}C_{F}}{\pi}\frac{3}{2}\left(-\frac{3}{4}+\frac{1}{2}\ln\left(\frac{Q^{2}}{x_{\star}P_{\perp}^{2}}\right)\right)$ (79)
C.4 2.4 Real diagrams without poles
Besides the virtual graphs in Fig. 2, there are contributions to the SIDIS jet cross-section coming from real diagrams without collinear divergence where the quark jet is measured and the antiquark and the gluon are integrated out. Since these diagrams have no collinear divergences, the jet definition does not play any role in the ”narrow jet” approximation $R\ll 1$. More formally, one can integrate over the full antiquark+gluon phase space, so that the resulting expression is in fact similar to the virtual corrections computed above. That is also why such diagrams actually contribute to the ”virtual” part of the NLO correction to the quark TMD ; we shall indeed observe that the quark jet $\bm{P}_{\perp}$ in these contributions comes from the quark TMD (“intrinsic $P_{\perp}$”) and not from the recoil associated with the gluon emission.
C.4.1 2.4.1 Real gluon emission from the unmeasured antiquark before the shock-wave
We start with the gluon emission by the antiquark before the SW both in the amplitude and in the complex conjugate amplitude. The respective contribution to the cross-section, denoted as $\overline{\rm R1}\times\overline{\rm R1}^{*}$, is UV singular because of the $1/\bm{r}_{zy}^{2}$ (with $\bm{r}_{zy}=\bm{z}_{\perp}-\bm{y}_{\perp}$) kernel in the integrand. To regularise this divergence, a UV regulator is subtracted from the expression for $\overline{\rm R1}\times\overline{\rm R1}^{*}$ and added to the pole term in Eq. (47) (cf. the discussion after Eq. (67)). After including this regulator, the diagram $\overline{\rm R1}\times\overline{\rm R1}^{*}$ reads
$\left.\frac{\mathrm{d}\sigma_{\rm CGC}^{\gamma_{\rm T}^{\star}+A\to q+X}}{\mathrm{d}^{2}\bm{b}_{\perp}\mathrm{d}^{2}\bm{P}_{\perp}\mathrm{d}z}\right|_{\overline{\rm R1}\times\overline{\rm R1}^{*}}$ $=\frac{\alpha_{\rm em}e_{f}^{2}N_{c}}{(2\pi)^{4}}\int\mathrm{d}^{2}\bm{r}_{xy}\mathrm{d}^{2}\bm{r}_{x^{\prime}y}\frac{\mathrm{d}^{2}\bm{r}_{zy}}{\bm{r}_{zy}^{2}}e^{-i\bm{P}_{\perp}\cdot\bm{r}_{xx^{\prime}}}\frac{\alpha_{s}}{2\pi^{2}}z^{2}(\bar{z}^{2}+(1-z)^{2})\left(z^{2}+(1-z)^{2}\right)$
×$\times\int_{0}^{1-z}\frac{\mathrm{d}z_{g}}{z_{g}}\left\{\left[Q^{2}K_{1}(QX_{R})K_{1}(QX_{R}^{\prime})\frac{\bm{R}_{\overline{\rm R}}\cdot\bm{R}_{\overline{\rm R}}^{\prime}}{X_{R}X_{R}^{\prime}}\,\Xi_{\rm NLO,4}(\bm{x}_{\perp},\bm{y}_{\perp},\bm{z}_{\perp},\bm{x}_{\perp}^{\prime})\right.\right.$
$\left.\left.-\frac{Q^{2}}{z(1-z)}K_{1}(\bar{Q}r_{xy})K_{1}(\bar{Q}r_{x^{\prime}y})e^{-\frac{\bm{r}_{zy}^{2}}{\bm{r}_{xx^{\prime}}^{2}e^{\gamma_{E}}}}\frac{\bm{r}_{xy}\cdot\bm{r}_{x^{\prime}y}}{r_{xy}r_{x^{\prime}y}}\,C_{F}\Xi_{\rm LO}(\bm{x}_{\perp},\bm{y}_{\perp},\bm{x}_{\perp}^{\prime})\right]\right\}_{z_{\star}}\,,$ (80)
with $\bm{R}_{\rm R}^{\prime}=\bm{r}_{x^{\prime}y}+\frac{z_{g}}{z+z_{g}}\bm{r}_{zx^{\prime}}$, $\bm{R}_{\overline{\rm R}}^{\prime}=-\bm{r}_{x^{\prime}y}+\frac{z_{g}}{1-z}\bm{r}_{zy}$, $X_{R}^{2}=z\bar{z}\bm{r}_{xy}^{2}+zz_{g}\bm{r}_{zx}^{2}+\bar{z}z_{g}\bm{r}_{zy}^{2}$ and $X_{R}^{\prime 2}=z\bar{z}\bm{r}_{x^{\prime}y}^{2}+zz_{g}\bm{r}_{zx^{\prime}}^{2}+\bar{z}z_{g}\bm{r}_{zy}^{2}$, $\bar{z}=1-z-z_{g}$. Once again, we do not show the instantaneous piece as it is power suppressed. The second term in the square bracket in Eq. (80) is meant to regulate this UV divergence, as just explained. The CGC correlator of this diagram is built from dipoles only,
$\Xi_{\rm NLO,4}(\bm{x}_{\perp},\bm{y}_{\perp},\bm{z}_{\perp},\bm{x}_{\perp}^{\prime})$ $\equiv\frac{N_{c}}{2}\left\langle D_{xx^{\prime}}-D_{xz}D_{zy}-D_{yz}D_{zx^{\prime}}+1\right\rangle-\frac{1}{2N_{c}}\left\langle D_{xx^{\prime}}-D_{xy}-D_{yx^{\prime}}+1\right\rangle$ (81)
Let now focus on the kinematic regime where the measured quark has $z\sim P_{\perp}^{2}/Q^{2}$ and $|\bm{z}_{\perp}-\bm{y}_{\perp}|\sim 1/Q\ll|\bm{x}_{\perp}-\bm{y}_{\perp}|\sim 1/P_{\perp}$. We have in mind that $|\bm{z}_{\perp}-\bm{y}_{\perp}|$, physically conjugated to the transverse momentum $\bm{k_{g\perp}}$ of the gluon, is controlled by sizes much smaller than $|\bm{x}_{\perp}-\bm{y}_{\perp}|=\mathcal{O}(1/P_{\perp})$, i.e. $\bm{k_{g\perp}}^{2}\gg P_{\perp}^{2}$. It implies that $\bm{R}_{\overline{\rm R}}\approx-\bm{r}_{xy}$. Note that the configuration where the measured quark has $1-z\sim P_{\perp}^{2}/Q^{2}$ is power suppressed since the integral over $z_{g}$ vanishes as $z\to 1$. In the limit $z\sim P_{\perp}^{2}/Q^{2}\ll 1$ and $r_{zy}\ll r_{xy}$, the (squared) transverse distance $X_{R}$ in the argument of the Bessel function reduces to
$X_{R}$ $=z(1-z-z_{g})\bm{r}_{xy}^{2}+zz_{g}\bm{r}_{zx}^{2}+(1-z-z_{g})z_{g}\bm{r}_{zy}^{2}$ (82)
$\approx z\bm{r}_{xy}^{2}+z_{g}(1-z_{g})\bm{r}_{zy}^{2}$ (83)
We also systematically expand the rational function in $z$ and $z_{g}$ by keeping the leading power piece in $z$ and we expand the color correlator $\Xi_{\rm NLO,4}$ around $\bm{z}_{\perp}=\bm{y}_{\perp}$ such that
$\Xi_{\rm NLO,4}(\bm{x}_{\perp},\bm{y}_{\perp},\bm{z}_{\perp},\bm{x}_{\perp}^{\prime})$ $\equiv\frac{N_{c}}{2}\left\langle D_{xx^{\prime}}-D_{xz}D_{zy}-D_{yz}D_{zx^{\prime}}+1\right\rangle-\frac{1}{2N_{c}}\left\langle D_{xx^{\prime}}-D_{xy}-D_{yx^{\prime}}+1\right\rangle\,,$
$=C_{F}\Xi_{\rm LO}(\bm{x}_{\perp},\bm{y}_{\perp},\bm{x}_{\perp}^{\prime})+\mathcal{O}(Q_{s}^{2}r_{zy}^{2})$ (84)
After all these approximations and going to momentum space, the cross-section (plus its complex conjugate) can be cast into the simpler form
$\left.\frac{\mathrm{d}\sigma^{\gamma_{\rm T}^{\star}+A\to q+X}}{\mathrm{d}^{2}\bm{b}_{\perp}\mathrm{d}^{2}\bm{P}_{\perp}}\right|_{\overline{\rm R1}}$ $=\frac{\alpha_{\rm em}e_{f}^{2}N_{c}}{(2\pi)}\frac{\alpha_{s}C_{F}}{\pi^{2}}\int\mathrm{d}^{2}\bm{q}_{\perp}\mathcal{D}(x,\bm{q}_{\perp})\int_{0}^{1}\mathrm{d}z\int_{x_{\star}P_{\perp}^{2}/Q^{2}}^{1}\frac{\mathrm{d}z_{g}}{z_{g}}\left[1+(1-z_{g})^{2}\right]$
×$\times\int\frac{\mathrm{d}^{2}\bm{l_{3\perp}}}{2\pi}\left[\mathcal{H}_{\overline{\rm R1}}^{ij*}\mathcal{H}_{\overline{\rm R1}}^{ij}-\mathcal{H}_{\overline{\rm R1},\rm uv}^{ij*}\mathcal{H}_{\overline{\rm R1},\rm uv}^{ij}\right]$ (85)
with the hard factor of diagram $\overline{\rm R1}$
$\mathcal{H}_{\overline{\rm R1}}^{ij}$ $\equiv\int\frac{\mathrm{d}^{2}\bm{r}_{zy}}{(2\pi)}\int\frac{\mathrm{d}^{2}\bm{r}_{xy}}{(2\pi)}e^{-i\bm{P}_{\perp}\cdot\bm{r}_{xy}}\left(e^{i\bm{q}_{\perp}\cdot\bm{r}_{xy}}-1\right)e^{i\bm{l_{3\perp}}\cdot\bm{r}_{zy}}\frac{\bm{r}_{zy}^{i}\bm{r}_{xy}^{j}}{\bm{r}_{zy}^{2}}\frac{\bar{Q}K_{1}\left(\bar{Q}\sqrt{\bm{r}_{xy}^{2}+\omega\bm{r}_{zy}^{2}}\right)}{\sqrt{\bm{r}_{xy}^{2}+\omega\bm{r}_{zy}^{2}}}$ (86)
$=-\frac{\bm{l_{3\perp}}^{i}(\bm{P}_{\perp}^{j}-\bm{q}_{\perp})^{j}}{\left[(\bm{P}_{\perp}-\bm{q}_{\perp})^{2}+\bar{Q}^{2}\right]\left[\bm{l_{3\perp}}^{2}+\omega((\bm{P}_{\perp}-\bm{q}_{\perp})^{2}+\bar{Q}^{2})\right]}+\frac{\bm{l_{3\perp}}^{i}\bm{P}_{\perp}^{j}}{\left[P_{\perp}^{2}+\bar{Q}^{2}\right]\left[\bm{l_{3\perp}}^{2}+\omega(P_{\perp}^{2}+\bar{Q}^{2})\right]}$ (87)
and its UV regulator,
$\mathcal{H}_{\overline{\rm R1},\rm uv}^{ij}$ $\equiv\int\frac{\mathrm{d}^{2}\bm{r}_{zy}}{(2\pi)}\int\frac{\mathrm{d}^{2}\bm{r}_{xy}}{(2\pi)}e^{-i\bm{P}_{\perp}\cdot\bm{r}_{xy}}\left(e^{i\bm{q}_{\perp}\cdot\bm{r}_{xy}}-1\right)e^{i\bm{l_{3\perp}}\cdot\bm{r}_{zy}}\frac{\bm{r}_{zy}^{i}\bm{r}_{xy}^{j}}{\bm{r}_{zy}^{2}}\frac{\bar{Q}K_{1}\left(\bar{Q}r_{xy}\right)\exp\left(-\frac{\bm{r}_{zy}^{2}}{4\xi}\right)}{r_{xy}}$ (88)
$=\left[\frac{(\bm{P}_{\perp}-\bm{q}_{\perp})^{j}}{(\bm{P}_{\perp}-\bm{q}_{\perp})^{2}+\bar{Q}^{2}}-\frac{\bm{P}_{\perp}^{j}}{P_{\perp}^{2}+\bar{Q}^{2}}\right]\frac{\bm{l_{3\perp}}^{i}}{\bm{l_{3\perp}}^{2}}\left(1-e^{-\xi\bm{l_{3\perp}}^{2}}\right)$ (89)
In principle, one choose the arbitrary parameter $\xi$ in the UV regulator to be $\xi=\bm{r}_{xx^{\prime}}^{2}e^{\gamma_{E}}/2$ to match the expression given by Eq. (80). This makes the calculation more difficult, although in the dilute limit, one is allowed to use $\xi=c_{0}/P_{\perp}^{2}$ instead (it gives identical finite pieces up to powers of $Q_{s}^{2}/P_{\perp}^{2}$ corrections). Integrating over the auxiliary variable $\bm{l_{3\perp}}$ gives
$\int\frac{\mathrm{d}^{2}\bm{l_{3\perp}}}{2\pi}\left[\mathcal{H}_{\overline{\rm R1}}^{ij*}\mathcal{H}_{\overline{\rm R1}}^{ij}-\mathcal{H}_{\overline{\rm R1},uv}^{ij*}\mathcal{H}_{\overline{\rm R1},uv}^{ij}\right]=\frac{(\bm{P}_{\perp}-\bm{q}_{\perp})^{2}}{2\left[(\bm{P}_{\perp}-\bm{q}_{\perp})^{2}+\bar{Q}^{2}\right]^{2}}\left[-1+\ln(c_{0})-\ln\left(\xi(\omega((\bm{P}_{\perp}-\bm{q}_{\perp})^{2}+\bar{Q}^{2}))\right)\right]$
$+\frac{P_{\perp}^{2}}{2\left[P_{\perp}^{2}+\bar{Q}^{2}\right]^{2}}\left[-1+\ln(c_{0})-\ln\left(\xi(\omega(P_{\perp}^{2}+\bar{Q}^{2}))\right)\right]-\frac{2\bm{P}_{\perp}\cdot(\bm{P}_{\perp}-\bm{q}_{\perp})}{2\left[(\bm{P}_{\perp}-\bm{q}_{\perp})^{2}+\bar{Q}^{2}\right]\left[P_{\perp}^{2}+\bar{Q}^{2}\right]}\Bigg{[}\ln(c_{0})-\ln(\omega\xi)$
$-\frac{\left[(\bm{P}_{\perp}-\bm{q}_{\perp})^{2}+\bar{Q}^{2}\right]\ln((\bm{P}_{\perp}-\bm{q}_{\perp})^{2}+\bar{Q}^{2})-\left[P_{\perp}^{2}+\bar{Q}^{2}\right]\ln(P_{\perp}^{2}+\bar{Q}^{2})}{(\bm{P}_{\perp}-\bm{q}_{\perp})^{2}-P_{\perp}^{2}}\Bigg{]}$ (90)
Finally, the take the dilute limite $P_{\perp}^{2}\gg q_{\perp}^{2}$ and we perform the integral over $z_{g}$ and $z$ (which is dominated by the endpoint $z=0$ so that the upper limit of the $z$ integration is irrelevant):
$\left.\frac{\mathrm{d}\sigma^{\gamma_{\rm T}^{\star}+A\to q+X}}{\mathrm{d}^{2}\bm{P}_{\perp}}\right|_{\overline{\rm R1}}=\left.\frac{\mathrm{d}\sigma^{\gamma_{\rm T}^{\star}+A\to q+X}}{\mathrm{d}^{2}\bm{P}_{\perp}}\right|_{\rm LO}\times\frac{\alpha_{s}C_{F}}{\pi}$ $\left[-\frac{3}{4}+\frac{\pi^{2}}{12}+\frac{3}{8}\ln\left(\frac{Q^{2}}{P_{\perp}^{2}}\right)-\frac{1}{2}\ln\left(\frac{Q^{2}}{x_{\star}P_{\perp}^{2}}\right)\ln\left(\frac{Q^{2}}{P_{\perp}^{2}}\right)+\frac{1}{4}\ln^{2}\left(\frac{Q^{2}}{x_{*}P_{\perp}^{2}}\right)\right.$
$\left.-\frac{5}{6}\left(-\frac{3}{4}+\ln\left(\frac{Q^{2}}{x_{\star}P_{\perp}^{2}}\right)\right)\right]$ (91)
[…]
C.4.2 2.4.2 Real gluon emission from the measured quark before the shock-wave
[…]
$\left.\frac{\mathrm{d}\sigma_{\rm CGC}^{\gamma_{\rm T}^{\star}+A\to q+X}}{\mathrm{d}^{2}\bm{b}_{\perp}\mathrm{d}^{2}\bm{P}_{\perp}\mathrm{d}z}\right|_{\rm R1\times R1^{*}}$ $=\frac{\alpha_{\rm em}e_{f}^{2}N_{c}}{(2\pi)^{4}}\int\mathrm{d}^{2}\bm{r}_{xy}\mathrm{d}^{2}\bm{r}_{x^{\prime}y}\mathrm{d}^{2}\bm{r}_{zy}e^{-i\bm{P}_{\perp}\cdot\bm{r}_{xx^{\prime}}}\frac{\alpha_{s}}{2\pi^{2}}Q^{2}\Xi_{\rm NLO,4}(\bm{x}_{\perp},\bm{y}_{\perp},\bm{z}_{\perp},\bm{x}_{\perp}^{\prime})$
×$\times\int_{0}^{1-z}\frac{\mathrm{d}z_{g}}{z_{g}}\left\{\frac{K_{1}(QX_{R})K_{1}(QX_{R}^{\prime})}{X_{R}X_{R}^{\prime}}\left[\bar{z}^{2}(z^{2}+(1-\bar{z})^{2})\left(\bar{z}^{2}+(1-\bar{z})^{2}\right)\frac{\bm{r}_{zx}\cdot\bm{r}_{zx^{\prime}}}{\bm{r}_{zx}^{2}\bm{r}_{zx^{\prime}}^{2}}\bm{R}_{\rm R}\cdot\bm{R}_{\rm R}^{\prime}\right.\right.$
$\left.\left.+\bar{z}^{2}z_{g}(1+z-\bar{z})(1-2\bar{z})\frac{\bm{r}_{zx}\times\bm{r}_{zx^{\prime}}}{\bm{r}_{zx}^{2}\bm{r}_{zx^{\prime}}^{2}}\bm{R}_{\rm R}\times\bm{R}_{\rm R}^{\prime}\right]\right\}_{z_{\star}}$ (92)
with $\bm{R}_{\rm R}=\bm{r}_{xy}+z_{g}/(z+z_{g})\bm{r}_{zx}$. Using similar techniques, the leading power piece of this diagram at the cross-section level is given by
$\left.\frac{\mathrm{d}\sigma^{\gamma_{\rm T}^{\star}+A\to q+X}}{\mathrm{d}^{2}\bm{b}_{\perp}\mathrm{d}^{2}\bm{P}_{\perp}}\right|_{\rm R1}$ $=\frac{\alpha_{\rm em}e_{f}^{2}N_{c}}{2\pi}\frac{\alpha_{s}C_{F}}{\pi^{2}}\int\mathrm{d}^{2}\bm{q}_{\perp}\mathcal{D}(x,\bm{q}_{\perp})\int_{0}^{1}\mathrm{d}z\int_{x_{\star}P_{\perp}^{2}/Q^{2}}^{1}\frac{\mathrm{d}z_{g}}{z_{g}}\left[z_{g}^{2}+(1-z_{g})^{2}\right]\int\frac{\mathrm{d}^{2}\bm{l_{3\perp}}}{(2\pi)}\mathcal{H}_{\rm R1}^{ij*}\mathcal{H}_{\rm R1}^{ij}$ (93)
Noting $\Delta=\sqrt{\omega}\bar{Q}$,
$\mathcal{H}_{\rm R1}^{ij}$ $\equiv\int\frac{\mathrm{d}^{2}\bm{r}_{zy}}{(2\pi)}\int\frac{\mathrm{d}^{2}\bm{r}_{xy}}{(2\pi)}e^{-i\bm{P}_{\perp}\cdot\bm{r}_{xy}}\left(e^{i\bm{q}_{\perp}\cdot\bm{r}_{xy}}-1\right)e^{i\bm{l_{3\perp}}\cdot\bm{r}_{zy}}\frac{\bm{r}_{zy}^{i}\bm{r}_{xy}^{j}}{\bm{r}_{xy}^{2}}\frac{\Delta K_{1}\left(\Delta\sqrt{\bm{r}_{zy}^{2}+\bm{r}_{xy}^{2}/\omega}\right)}{\sqrt{\bm{r}_{zy}^{2}+\bm{r}_{xy}^{2}/\omega}}$ (94)
$=-\frac{\bm{l_{3\perp}}^{i}(\bm{P}_{\perp}^{j}-\bm{q}_{\perp})^{j}}{\left[\bm{l_{3\perp}}^{2}/\omega+\bar{Q}^{2}\right]\left[\bm{l_{3\perp}}^{2}+\omega((\bm{P}_{\perp}-\bm{q}_{\perp})^{2}+\bar{Q}^{2})\right]}+\frac{\bm{l_{3\perp}}^{i}\bm{P}_{\perp}^{j}}{\left[\bm{l_{3\perp}}^{2}/\omega+\bar{Q}^{2}\right]\left[\bm{l_{3\perp}}^{2}+\omega(P_{\perp}^{2}+\bar{Q}^{2})\right]}$ (95)
The integral over $\bm{l_{3\perp}}$ is straightforward, and one gets
$\int\frac{\mathrm{d}^{2}\bm{l_{3\perp}}}{(2\pi)}\mathcal{H}_{\rm R1}^{ij*}\mathcal{H}_{\rm R1}^{ij}$ $=\frac{-2(\bm{P}_{\perp}-\bm{q}_{\perp})^{2}+((\bm{P}_{\perp}-\bm{q}_{\perp})^{2}+2\bar{Q}^{2})\ln\left(1+\frac{(\bm{P}_{\perp}-\bm{q}_{\perp})^{2}}{\bar{Q}^{2}}\right)}{2(\bm{P}_{\perp}-\bm{q}_{\perp})^{4}}+\frac{-2P_{\perp}^{2}+(P_{\perp}^{2}+2\bar{Q}^{2})\ln\left(1+\frac{P_{\perp}^{2}}{\bar{Q}^{2}}\right)}{2\bm{P}_{\perp}^{4}}$
$-\frac{(\bm{P}_{\perp}-\bm{q}_{\perp})\cdot\bm{P}_{\perp}}{(\bm{P}_{\perp}-\bm{q}_{\perp})^{4}\bm{P}_{\perp}^{4}((\bm{P}_{\perp}-\bm{q}_{\perp})^{2}-P_{\perp}^{2})}\Bigg{[}(\bm{P}_{\perp}-\bm{q}_{\perp})^{2}\bm{P}_{\perp}^{4}-P_{\perp}^{2}(\bm{P}_{\perp}-\bm{q}_{\perp})^{4}$
$-\bm{P}_{\perp}^{4}((\bm{P}_{\perp}-\bm{q}_{\perp})^{2}+\bar{Q}^{2})\ln\left(1+\frac{(\bm{P}_{\perp}-\bm{q}_{\perp})^{2}}{\bar{Q}^{2}}\right)+(\bm{P}_{\perp}-\bm{q}_{\perp})^{4}(P_{\perp}^{2}+\bar{Q}^{2})\ln\left(1+\frac{P_{\perp}^{2}}{\bar{Q}^{2}}\right)\Bigg{]}$ (96)
This diagram does not generate Sudakov double logarithm, but it contributes to the Sudakov single logarithm. Expanding for $P_{\perp}\gg q_{\perp}$, we get
$\left.\frac{\mathrm{d}\sigma^{\gamma_{\rm T}^{\star}+A\to q+X}}{\mathrm{d}^{2}\bm{P}_{\perp}}\right|_{\rm R1}$ $=\left.\frac{\mathrm{d}\sigma^{\gamma_{\rm T}^{\star}+A\to q+X}}{\mathrm{d}^{2}\bm{P}_{\perp}}\right|_{\rm LO}\times\frac{(-\alpha_{s})C_{F}}{8\pi}\left[1-\ln\left(\frac{Q^{2}}{x_{\star}P_{\perp}^{2}}\right)\right]$ (97)
C.4.3 2.4.3 Interferences between real gluon emission before the shock-wave from the measured quark and the unmeasured antiquark
[…]
$\left.\frac{\mathrm{d}\sigma_{\rm CGC}^{\gamma_{\rm T}^{\star}+A\to q+X}}{\mathrm{d}^{2}\bm{b}_{\perp}\mathrm{d}^{2}\bm{P}_{\perp}\mathrm{d}z}\right|_{\rm R1\times\overline{R1}^{*}}$ $=\frac{\alpha_{\rm em}e_{f}^{2}N_{c}}{(2\pi)^{4}}\int\mathrm{d}^{2}\bm{r}_{xy}\mathrm{d}^{2}\bm{r}_{x^{\prime}y}\mathrm{d}^{2}\bm{r}_{zy}e^{-i\bm{P}_{\perp}\cdot\bm{r}_{xx^{\prime}}}\frac{\alpha_{s}}{2\pi^{2}}2zQ^{2}\Xi_{\rm NLO,4}(\bm{x}_{\perp},\bm{y}_{\perp},\bm{z}_{\perp},\bm{x}_{\perp}^{\prime})$
×$\times\int_{0}^{1-z}\frac{\mathrm{d}z_{g}}{z_{g}}\left\{\frac{K_{1}(QX_{R})K_{1}(QX_{R}^{\prime})}{X_{R}X_{R}^{\prime}}\left[\bar{z}(z(1-z)+\bar{z}(1-\bar{z}))(z(1-\bar{z})+\bar{z}(1-z))\frac{\bm{r}_{zx}\cdot\bm{r}_{zy}}{\bm{r}_{zx}^{2}\bm{r}_{zy}^{2}}\bm{R}_{\rm R}\cdot\bm{R}_{\overline{\rm R}}^{\prime}\right.\right.$
$\left.\left.-\bar{z}z_{g}(z-\bar{z})^{2}\frac{\bm{r}_{zx}\times\bm{r}_{zy}}{\bm{r}_{zx}^{2}\bm{r}_{zy}^{2}}\bm{R}_{\rm R}\times\bm{R}_{\overline{\rm R}}^{\prime}\right]\right\}_{z_{\star}}$ (98)
where we recall that $\bar{z}=1-z-z_{g}$. In the limit $r_{zy}\ll r_{xy}$ and $z\ll 1$, the two terms nicely combine to give
$2z^{2}(1-z_{g})^{3}z_{g}\frac{(\bm{r}_{xy}\cdot\bm{r}_{zy})(\bm{r}_{zy}\cdot\bm{r}_{x^{\prime}y})-(\bm{r}_{xy}\times\bm{r}_{zy})(\bm{r}_{zy}\times\bm{r}_{x^{\prime}y})}{\bm{r}_{xy}^{2}\bm{r}_{zy}^{2}}$ $=2z^{2}(1-z_{g})^{3}z_{g}\frac{\bm{r}_{xy}\cdot\bm{r}_{x^{\prime}y}}{\bm{r}_{xy}^{2}}$ (99)
Including the complex conjugate diagram and using the same approximations as above, we find
$\left.\frac{\mathrm{d}\sigma^{\gamma_{\rm T}^{\star}+A\to q+X}}{\mathrm{d}^{2}\bm{b}_{\perp}\mathrm{d}^{2}\bm{P}_{\perp}}\right|_{\rm int}$ $=\frac{\alpha_{\rm em}e_{f}^{2}N_{c}}{2\pi}\frac{2\alpha_{s}C_{F}}{\pi^{2}}\int\mathrm{d}^{2}\bm{q}_{\perp}\mathcal{D}(x,\bm{q}_{\perp})\int_{0}^{1}\frac{\mathrm{d}z}{z}\int_{x_{\star}P_{\perp}^{2}/Q^{2}}^{1}\frac{\mathrm{d}z_{g}}{z_{g}}(1-z_{g})^{3}z_{g}\int\frac{\mathrm{d}^{2}\bm{l_{3\perp}}}{(2\pi)}\mathcal{H}_{\rm int,a}^{i}\mathcal{H}_{\rm int,b}^{i*}$ (100)
with the hard factor defined as
$\mathcal{H}^{i}_{\rm int,a}$ $\equiv\int\frac{\mathrm{d}^{2}\bm{r}_{zy}}{(2\pi)}\int\frac{\mathrm{d}^{2}\bm{r}_{xy}}{(2\pi)}e^{-i\bm{P}_{\perp}\cdot\bm{r}_{xy}}\left(e^{i\bm{q}_{\perp}\cdot\bm{r}_{xy}}-1\right)e^{i\bm{l_{3\perp}}\cdot\bm{r}_{zy}}\frac{\bm{r}_{xy}^{i}}{\bm{r}_{xy}^{2}}\frac{\bar{Q}K_{1}\left(\bar{Q}\sqrt{\bm{r}_{xy}^{2}+\omega\bm{r}_{zy}^{2}}\right)}{\sqrt{\bm{r}_{xy}^{2}+\omega\bm{r}_{xy}^{2}}}$ (101)
$=\frac{(-i)}{2\omega}\left[\frac{\bm{P}_{\perp}^{i}-\bm{q}_{\perp}^{i}}{(\bm{P}_{\perp}-\bm{q}_{\perp})^{2}}\ln\left(1+\frac{(\bm{P}_{\perp}-\bm{q}_{\perp})^{2}}{\bar{Q}^{2}+\bm{l_{3\perp}}^{2}/\omega}\right)-\frac{\bm{P}_{\perp}^{i}}{P_{\perp}^{2}}\ln\left(1+\frac{P_{\perp}^{2}}{\bar{Q}^{2}+\bm{l_{3\perp}}^{2}/\omega}\right)\right]$ (102)
$\mathcal{H}^{i}_{\rm int,b}$ $\equiv\int\frac{\mathrm{d}^{2}\bm{r}_{zy}}{(2\pi)}\int\frac{\mathrm{d}^{2}\bm{r}_{xy}}{(2\pi)}e^{-i\bm{P}_{\perp}\cdot\bm{r}_{xy}}\left(e^{i\bm{q}_{\perp}\cdot\bm{r}_{xy}}-1\right)e^{i\bm{l_{3\perp}}\cdot\bm{r}_{zy}}\bm{r}_{xy}^{i}\frac{\bar{Q}K_{1}\left(\bar{Q}\sqrt{\bm{r}_{xy}^{2}+\omega\bm{r}_{zy}^{2}}\right)}{\sqrt{\bm{r}_{xy}^{2}+\omega\bm{r}_{xy}^{2}}}$ (103)
$=\frac{-2i\omega(\bm{P}_{\perp}-\bm{q}_{\perp})^{i}}{\left[\bm{l_{3\perp}}^{2}+\omega((\bm{P}_{\perp}-\bm{q}_{\perp})^{2}+\bar{Q}^{2}\right]^{2}}+\frac{2i\omega\bm{P}_{\perp}^{i}}{\left[\bm{l_{3\perp}}^{2}+\omega(P_{\perp}^{2}+\bar{Q}^{2}\right]^{2}}$ (104)
The $\bm{l_{3\perp}}$ integration is again straightforward since
$\int\frac{\mathrm{d}^{2}\bm{l_{3\perp}}}{(2\pi)}\mathcal{H}_{\rm int,a}^{i}\mathcal{H}_{\rm int,b}^{i*}=\frac{P_{\perp}^{2}-\bar{Q}^{2}\ln\left(1+\frac{P_{\perp}^{2}}{\bar{Q}^{2}}\right)}{2\omega P_{\perp}^{2}(P_{\perp}^{2}+\bar{Q}^{2})}+\frac{(\bm{P}_{\perp}-\bm{q}_{\perp})^{2}-\bar{Q}^{2}\ln\left(1+\frac{(\bm{P}_{\perp}-\bm{q}_{\perp})^{2}}{\bar{Q}^{2}}\right)}{2\omega(\bm{P}_{\perp}-\bm{q}_{\perp})^{2}((\bm{P}_{\perp}-\bm{q}_{\perp})^{2}+\bar{Q}^{2})}$
$+\frac{\bm{P}_{\perp}\cdot(\bm{P}_{\perp}-\bm{q}_{\perp})\bar{Q}^{2}}{2\omega(\bm{P}_{\perp}-\bm{q}_{\perp})^{2}P_{\perp}^{2}\left[P_{\perp}^{2}+\bar{Q}^{2}\right]}\ln\left(1+\frac{(\bm{P}_{\perp}-\bm{q}_{\perp})^{2}}{\bar{Q}^{2}}\right)+\frac{\bm{P}_{\perp}\cdot(\bm{P}_{\perp}-\bm{q}_{\perp})}{2\omega P_{\perp}^{2}\left[P_{\perp}^{2}-(\bm{P}_{\perp}-\bm{q}_{\perp})^{2})\right]}\ln\left(\frac{(\bm{P}_{\perp}-\bm{q}_{\perp})^{2}+\bar{Q}^{2}}{P_{\perp}^{2}+\bar{Q}^{2}}\right)$
$+\frac{\bm{P}_{\perp}\cdot(\bm{P}_{\perp}-\bm{q}_{\perp})\bar{Q}^{2}}{2\omega(\bm{P}_{\perp}-\bm{q}_{\perp})^{2}P_{\perp}^{2}\left[(\bm{P}_{\perp}-\bm{q}_{\perp})^{2}+\bar{Q}^{2}\right]}\ln\left(1+\frac{P_{\perp}^{2}}{\bar{Q}^{2}}\right)+\frac{\bm{P}_{\perp}\cdot(\bm{P}_{\perp}-\bm{q}_{\perp})}{2\omega(\bm{P}_{\perp}-\bm{q}_{\perp})^{2}\left[P_{\perp}^{2}-(\bm{P}_{\perp}-\bm{q}_{\perp})^{2})\right]}\ln\left(\frac{(\bm{P}_{\perp}-\bm{q}_{\perp})^{2}+\bar{Q}^{2}}{P_{\perp}^{2}+\bar{Q}^{2}}\right)$ (105)
Note that the additional $1/z$ factor in the $z_{g}$ integration combines with the $1/\omega=z/(z_{g}(1-z_{g})$ to give a leading power result. In particular, in the dilute limit,
$\left.\frac{\mathrm{d}\sigma^{\gamma_{\rm T}^{\star}+A\to q+X}}{\mathrm{d}^{2}\bm{P}_{\perp}}\right|_{\rm int}$ $=\left.\frac{\mathrm{d}\sigma^{\gamma_{\rm T}^{\star}+A\to q+X}}{\mathrm{d}^{2}\bm{P}_{\perp}}\right|_{\rm LO}\times\frac{\alpha_{s}C_{F}}{4\pi}\left[-\frac{3}{2}+\ln\left(\frac{Q^{2}}{x_{\star}P_{\perp}^{2}}\right)\right]$ (106)
C.5 2.5 Summary
To sum up this lengthy calculation, we combine together all our intermediate results for the NLO diagrams which contribute at leading twist, namely Eq. (56)-(57), Eq. (70), Eq. (79), Eq. (91), Eq. (97) and Eq. (106). For the clustering condition A, we get
$\left.\frac{\mathrm{d}\sigma^{\gamma^{*}_{\rm T}+A\to j+X}}{\mathrm{d}^{2}\bm{P}_{\perp}}\right|_{\rm NLO,A-def}=\left.\frac{\mathrm{d}\sigma^{\gamma^{*}_{\rm T}+A\to j+X}}{\mathrm{d}^{2}\bm{P}_{\perp}}\right|_{\rm LO}$$\times\frac{\alpha_{s}C_{F}}{\pi}\left[-\frac{3}{4}\ln^{2}\left(\frac{Q^{2}}{P_{\perp}^{2}}\right)+\left(\frac{3}{4}-\ln(R)\right)\ln\left(\frac{Q^{2}}{P_{\perp}^{2}}\right)\right.$
$\left.-\frac{3}{2}\ln(2R)+\frac{17}{4}-\frac{5\pi^{2}}{12}+\frac{3}{4}\ln^{2}(x_{\star})+\frac{3}{8}\ln(x_{\star})+\mathcal{O}(R^{2})\right]\,.$ (107)
while for the clustering condition B,
$\left.\frac{\mathrm{d}\sigma^{\gamma^{*}_{\rm T}+A\to j+X}}{\mathrm{d}^{2}\bm{P}_{\perp}}\right|_{\rm NLO,B-def}=\left.\frac{\mathrm{d}\sigma^{\gamma^{*}_{\rm T}+A\to j+X}}{\mathrm{d}^{2}\bm{P}_{\perp}}\right|_{\rm LO}$$\times\frac{\alpha_{s}C_{F}}{\pi}\left[-\frac{1}{4}\ln^{2}\left(\frac{Q^{2}}{P_{\perp}^{2}}\right)+\left(\frac{3(1-\beta/2)}{4}+\ln(R)\right)\ln\left(\frac{Q^{2}}{P_{\perp}^{2}}\right)\right.$
$\left.-\frac{3}{2}\ln(R)+\frac{11}{4}-\frac{3\pi^{2}}{4}+\frac{3}{4}\ln^{2}(x_{\star})+\frac{3}{8}\ln(x_{\star})+\mathcal{O}(R^{2})\right]\,.$ (108)
An important consistency check of these results is the fact that they do not involve mixed double logarithms of the type shown in Eq. (71). This is important since it guarantees that the small-$x$ and the Sudakov resummations (equivalently, the B-JIMWLK and the CSS evolutions) can be performed simultaneously without overlap in phase space. In particular, this property is consistent with the fact that the CSS evolution is local in $x$. Moreover, the $x_{\star}$–dependent remainder is identical for both jet definitions; this shows that small-$x$ evolution is not sensitive to the final state collinear dynamic. In both cases, we recognise the same term of the form $-(3/2)\ln(R)$ coming from final-state, hard and collinear, gluon splittings.
Let us now turn to the differences between the two classes of jet definitions, $A$ and $B$, as manifest from the inspection of the Sudakov logarithms in Eqs. (107) and (108). Clearly, the main difference refers to the coefficient of the Sudakov double logarithm. With the clustering condition $B$, this coefficient is equal to $-\alpha_{s}C_{F}/(4\pi)$ (like in Eq. (12) in the Letter) and is consistent with TMD factorisation, as also explained in the Letter: (i) it has the right value to cancel the (single) real Sudakov log in Eq. (11) of the Letter after integrating over $P_{\perp}$, and (ii) it is furthermore consistent with expectations from the CSS formalism Collins and Soper (1981, 1982); Collins et al. (1985); Collins (2023). On the contrary, the respective coefficient predicted by the clustering condition $B$, that is, $-3\alpha_{s}C_{F}/(4\pi)$, not only is inconsistent with the general CSS results aforementioned, but also does not fulfil the condition of real–virtual cancellation after integrating over $P_{\perp}$; indeed, with this jet definition, the real Sudakov logarithm is found to be twice as large as that shown in the second line of Eq. (11), hence it cannot cancel against a virtual double logarithm which is three times larger than that in Eq. (108).
As furthermore explained in the Letter, the coefficient of the Sudakov single logarithm is important too to ensure TMD factorisation, namely it needs to match the quark anomalous dimension $\Gamma_{q}=\frac{3\alpha_{s}C_{F}}{4\pi}$ and by the same token to provide the correct virtual piece of the splitting function $\mathcal{P}_{qq}(\xi)$ (recall the discussion after Eq. (13) in the Letter). By inspection of Eq. (108), it is clear that these conditions are satisfied if and only we choose $\beta=0$. We recall that in the narrow jet approximation ($R\ll 1$), $\beta=1$ corresponds to the inclusive $k_{t}$ algorithm designed in Catani et al. (1992), while $\beta=2$ corresponds the inclusive generalised-$k_{t}$ algorithm for $e^{+}e^{-}$ implemented in the Breit frame. Yet, in both cases, these widely used jet definitions fail to correctly account for the single Sudakov logarithm. In particular, with $\beta=2$, the respective coefficient simply vanishes!
From a physical point of view, the case $\beta=2$ can be understood as follows. Consider the situation where, in the dipole frame, the tagged quark is slow, with $z_{1}\ll 1$. That is, the fermion struck by the photon in the Breit frame is the (unmeasured) antiquark. Still in the Breit frame, the tagged jet propagates in the fragmentation region of the nuclear target. If this jet is reconstructed using the inclusive generalised-$k_{t}$ algorithm for $e^{+}e^{-}$ (i.e. Eq. (32) with $\beta=2$) and if the gluon is close in angle to the tagged quark with $z_{1}\ll 1$, then both particles are recombined into the same jet with $z=z_{g}+z_{1}\ll 1$. In such a case, the gluon emission does not modify the structure of the final state, so there is a complete cancellation of the single Sudakov logarithms between real (in-cone) and virtual emissions, cf. Eq. (108) with $\beta=2$. On the other hand, when using the distance measure given by Eq. (5) in the Letter with $p=1$ (or Eq. (32) with $\beta=0$), the gluon and the quark jet with $z_{g},z_{1}\ll 1$ are strongly penalised by the clustering criterion: the left hand side of Eq. (32) becomes large for $z_{1}z_{g}\ll 1$, while its r.h.s. is equal to 1 for $\beta=0$. So, these very soft quark-gluon pairs are typically not recombined inside the same jet by our new clustering condition, meaning that the virtual contributions to the single Sudakov logs remain uncompensated, as indeed visible in Eq. (108) when $\beta=0$.
To summarise, the physical jet definition for jets in SIDIS — the one to be consistent with TMD factorisation at NLO — is the one that we propose in Eq. (5) in the Letter with the choice $p=1$, or, equivalently, that shown in Eq. (32) of this Supplemental Material, with the choice $\beta=0$. For this definition, the result that we have just derived for the NLO virtual corrections, i.e. Eq. (108) with $\beta=0$, coincides indeed with that shown in Eq. (12) in the Letter.
Appendix D 3 General Discussion on TMD Factorisation for SIDIS with the Jet Algorithm from Eq. (5)
In this section, we shall present an alternative argument for TMD factorisation, which is formulated in the more traditional framework of the target picture — that is, it follows the dynamics of the quark struck by the virtual photon. This argument neglects the saturation effects, hence it applies for moderate values of $x$, and also at small $x$ but sufficiently large values for the jet transverse momentum $P_{\perp}$, within the range $Q^{2}\gg P_{\perp}^{2}\gg Q_{s}^{2}(x)$. In this framework, one can rely on well established techniques to compute the NLO corrections to both the SIDIS cross-section and the quark TMD (starting with its operator definition). Previous calculations of SIDIS have focused on the measurement of a hadron (see for instance, the QCD factorisation and one-loop calculations in Refs. Ji et al. (2005); Collins (2023); Sun and Yuan (2013)). Here, we shall rather consider the production of a jet, with the purpose of demonstrating that — in this context too — the proper definition of the jet is essential in order to achieve TMD factorisation. Namely, we will show that the NLO corrections to the jet cross-section match the expected DGLAP+CSS evolution of the quark TMD if and only if one uses the jet definition given by Eq. (5) in the main text (or, equivalently, Eq. (32) with $\beta=0$ in this Supplemental Material).
Consider jet production in SIDIS, that is, the process $e+A\to e^{\prime}+\textrm{jet}(P_{\perp})+X$, in the (hard) TMD kinematics at $Q^{2}\gg P_{\perp}^{2}\gg Q_{s}^{2}(x)$. We shall parametrise the corresponding cross-section as follows,
$\frac{\mathrm{d}^{4}\sigma^{e+A\to e^{\prime}+j+X}}{\mathrm{d}x\mathrm{d}y\mathrm{d}^{2}\bm{P}_{\perp}}$ $=$ $\sigma_{0}^{\rm(DIS)}\,xF_{UU}(x,\bm{P}_{\perp},Q)\ ,$ (109)
where $x\equiv x_{\rm Bj}$, $\bm{P}_{\perp}$ is the transverse momentum of the produced jet with respect to the virtual photon direction, and (with the usual DIS kinematic variables $y$, $x_{\rm Bj}$, $Q^{2}$, and $W^{2}$, and $S_{ep}=(\ell_{e}+P_{N})^{2}$)
$\sigma_{0}^{\rm(DIS)}=\,\frac{4\pi\alpha^{2}_{\rm em}e_{f}^{2}S_{ep}}{Q^{4}}\,(1-y+y^{2}/2)\,.$ (110)
Note that with respect to Eq. (1) in the main text, we include here the leptonic part of the cross-section. $F_{UU}$ represents the spin-averaged structure function and. At LO, it is given by the (LO) quark TMD, cf. Eq. (1) in the Letter: $xF_{UU}^{(0)}=x\mathcal{F}^{(0)}_{q}(x,\bm{P}_{\perp})$.
As announced, we consider this process in the target picture and assume that the LO approximation corresponds to a valence quark, which is collinear with the nuclear target, $k_{q}^{\mu}=xP_{N}^{\mu}$, so in particular it has zero intrinsic transverse momentum:
$xF_{UU}^{(0)}(x,\bm{P}_{\perp})\,=\,\delta^{(2)}(\bm{P}_{\perp})\,xf_{q}^{(0))}(x),$ (111)
where $xf_{q}^{(0)}(x)$ represents the integrated quark distribution in the parton model and is scale independent.
For the purposes of the TMD factorisation, and notably in order to deal with low values for the transverse momentum, it is more convenient to work in the impact parameter representation, as obtained after a Fourier transform:
$xF_{UU}(x,\bm{P}_{\perp},Q)$ $=$ $\int\frac{\mathrm{d}^{2}\bm{b}_{\perp}}{(2\pi)^{2}}e^{i\bm{P}_{\perp}\cdot\bm{b}_{\perp}}x\widetilde{F}_{UU}(x,\bm{b}_{\perp},Q)\ .$ (112)
The Eq. (111) implies, at tree-level, $x\widetilde{F}_{UU}(x,\bm{b}_{\perp},Q)$ is independent upon both $\bm{b}_{\perp}$ and $Q^{2}$:
$x\tilde{F}^{(0)}_{UU}(x,\bm{b}_{\perp},Q)=xf_{q}^{(0)}(x).$ (113)
Figure 3: Feynman diagrams for the calculation of the NLO corrections to the SIDIS jet cross-section.
Before we proceed, let us emphasise an interesting difference w.r.t. the previous calculation using the dipole picture: in the present case, the measured jet can only be generated by the struck quark, so it necessarily carries a large fraction $z_{1}\equiv(k_{1}\cdot P_{N})/(q\cdot P_{N})\simeq 1$ of the longitudinal momentum of the virtual photon. This has two important implications: (i) in the final state, there is no jet in the fragmentation region of the nuclear target, and (ii) the jet cross-section has no sensitivity to the parameter $\beta$ which enters the clustering condition $B$, Eq. (32). Indeed, the jet has a longitudinal fraction $z\gtrsim z_{1}\simeq 1$ and then the $\beta$–dependence of Eq. (32) becomes irrelevant at NLO (altough it still matters for higher order calculations, starting at two loop order). This discussion also shows that, unlike in the previous scenario at small $x$ (where the measured jet can be sourced by either fermion from a sea quark-antiquark pair), the present calculation at moderate $x$ cannot discriminate between different jet definitions from the class represented by Eq. (32). With this in mind, let us return to the calculation of the jet cross-section for the case where at tree-level we have just a valence quark.
Clearly, the LO (or tree-level) approximation corresponds to the familiar “handbag” diagram for the absorption of the virtual photon by a valence quark. The NLO corrections are associated with the three amplitude graphs depicted in Fig. 3: real gluon emission in the initial state, or in the final state (i.e. prior and respectively after the photon absorption) and the virtual correction to the quark-photon vertex. (As before, we work in dimensional regularisation, so the quark self-energies are identically zero.) The virtual graph in Fig. 3 gives the following contribution in the $\overline{\rm MS}$ scheme:
${\left.\tilde{F}_{UU}(x,\bm{b}_{\perp},Q)\right|_{\cal V}}=\frac{\alpha_{s}}{2\pi}C_{F}\left(\frac{\mu^{2}}{Q^{2}}\right)^{\epsilon}\left\{-\frac{2}{\epsilon^{2}}-\frac{3}{\epsilon}-8\right\}\,.$ (114)
This does not modify the transverse momentum dependence, leading to a constant contribution in $b_{\perp}$-space.
Another contribution which brings no momentum dependence and hence can be effectively seen as a virtual correction, is the emission of a real gluon in the final state (cf. the middle graph in Fig. 3), but such the gluon and the outgoing quark are clustered inside the same jet. This is known as “the jet contribution” and is similar to Eq. (48) in the small $x$ calculation of section 2. As detailed in the previous section, the calculation of this contribution is strongly sensitive to the jet definition. We compute this contribution by using the jet algorithm in Eq. (5) of the main text and thus find
$J_{q}$ $=$ $\frac{\alpha_{s}}{2\pi}C_{F}\left(\frac{\mu^{2}}{Q^{2}}\right)^{\epsilon}\left\{\frac{1}{\epsilon^{2}}+\frac{1}{\epsilon}\ln\frac{1}{R^{2}}+\frac{3}{2}\frac{1}{\epsilon}+\frac{1}{2}\ln^{2}\frac{1}{R^{2}}+\frac{3}{2}\ln\frac{1}{R^{2}}+\frac{13}{2}-\frac{2}{3}\pi^{2}\right\}.$ (115)
Consider now the real gluon emissions where the gluon jet and the quark jet are well separated from each other (in the sense of our jet definition, of course). This emissions provide the transverse momentum $\bm{P}_{\perp}$ of the final jet via their recoil. For more clarity, we denote the 4-momenta of the incoming quark, the outgoing quark and the gluon as $k_{q}^{\mu}$, $k_{1}^{\mu}$, and $k_{g}^{\mu}$, respectively. Clearly, $\bm{k}_{q\perp}=0$ and $\bm{k}_{1\perp}=\bm{P}_{\perp}=-\bm{k}_{g\perp}$. To compute gluon radiation, it is convenient to chose the physical polarisation of the radiated gluon along the direction of the incoming hadron: $\varepsilon(k_{g})\cdot k_{q}=0$. This is convenient since in this gauge there is no soft radiation by the incoming quark, but only by the outgoing one. If we chose the hadron to be a left mover (as throughout this paper), then $k_{q}$ has only a minus component, $k_{q}^{\mu}=\delta^{\mu-}xP_{N}^{-}$, and the gauge condition becomes $A^{+}=0$ (the LC gauge of the virtual photon projectile).
Consider first soft gluon emission by the final quark. In the eikonal approximation, the respective contribution to the cross-section factorises and is proportional to
$g^{2}C_{F}\,\frac{k_{1}^{\mu}}{2k_{g}\cdot k_{1}-i\epsilon}\,\frac{k_{1}^{\nu}}{2k_{g}\dot{k}_{1}+i\epsilon}\,\left(-g^{\mu\nu}+\frac{k_{g}^{\mu}k_{q}^{\nu}+k_{g}^{\nu}k_{q}^{\mu}}{k_{g}\cdot k_{q}}\right)=g^{2}C_{F}S_{g}(k_{q},k_{1},k_{g}),$ (116)
with the notation
$S_{g}(k_{q},k_{1},k_{g})\equiv\,\frac{2k_{q}\cdot k_{1}}{(k_{q}\cdot k_{g})(k_{1}\cdot k_{g})}\,\simeq\,\frac{4}{k_{g\perp}^{2}}\,,$ (117)
where the final, approximate, equality holds since $k_{q}\cdot k_{1}=k_{q}^{-}k_{1}^{+}=xz_{1}P_{N}^{-}q^{+}$, $k_{q}\cdot k_{g}=xz_{g}P_{N}^{-}q^{+}$, and
$2k_{1}\cdot k_{g}=\frac{z_{g}}{z_{1}}k_{1\perp}^{2}+\frac{z_{1}}{z_{g}}k_{g\perp}^{2}-2\bm{k}_{1\perp}\cdot\bm{k}_{g\perp}\,=\,\frac{z_{1}}{z_{g}}\left(\bm{k}_{g\perp}-\frac{z_{g}}{z_{1}}\bm{k}_{1\perp}\right)^{2}\simeq\,\frac{z_{1}}{z_{g}}\,k_{g\perp}^{2}$ (118)
where we have used $z_{g}\ll z_{1}\simeq 1$ and $k_{1\perp}=k_{g\perp}=P_{\perp}$. The final result in Eq. (116) is gauge independent and can alternatively be interpreted in the covariant gauge as the “antenna pattern” produced via soft radiation from both the incoming and the outgoing quarks, including their interference.
To obtain the contribution of the final-state radiation to the jet cross-section, we need to integrate out the phase space of $k_{g}$, while keeping its transverse momentum fixed to $-\bm{P}_{\perp}$ and excluding gluon radiation inside the jet:
$\mathcal{S}_{\rm out}$ $\equiv 4\pi\alpha_{s}C_{F}\int\frac{\mathrm{d}k_{g}^{+}\mathrm{d}^{2}\bm{k_{g\perp}}}{2(2\pi)^{3}k_{g}^{+}}\,\delta^{(2)}(\bm{k_{g\perp}}+\bm{P}_{\perp})\,S_{g}(k_{q},k_{1},k_{g})\,\Theta\left((k_{1}+k_{g})^{2}>Q^{2}z_{g}R^{2}\right)$
$\simeq\frac{\alpha_{s}C_{F}}{\pi^{2}}\frac{1}{P_{\perp}^{2}}\int\frac{\mathrm{d}z_{g}}{z_{g}}\,\Theta\left(P_{\perp}^{2}>Q^{2}z_{g}^{2}R^{2}\right)\ ,$ (119)
where we have applied the jet algorithm of Eq. (5) with $p=1$. Recalling the lower limit $z_{g}\gg P_{\perp}^{2}/Q^{2}$ on $z_{g}$ (the boundary of the phase-space for high-energy evolution), we obtain
$\mathcal{S}_{\rm out}=\frac{\alpha_{s}C_{F}}{2\pi^{2}}\frac{1}{P_{\perp}^{2}}\ln\left(\frac{Q^{2}}{R^{2}P_{\perp}^{2}}\right).$ (120)
This is recognised as the real Sudakov logarithm previously obtained in the dipole picture, cf. the second line in Eq. (11) of the Letter. This result is clearly sensitive to the value of the parameter $p$ in the jet definition, Eq. (5) of the main text: for a different value of $p$, like $p=0$, the coefficient in front of the Sudakov logarithm would be different, which in turn would entail a mismatch between the real and the virtual NLO corrections, to be shortly computed.
To complete the calculation of the real NLO contribution to $F_{UU}$ in the approximation of interest, we need to add the contribution of the collinear gluon radiation from the incoming quark. A standard calculation using dimensional regularisation yields
$xF_{UU}(x,\bm{P}_{\perp},Q)\Big{|}_{\mathcal{R}}$ $=$ $\frac{\alpha_{s}C_{F}}{2\pi^{2}}\frac{1}{P_{\perp}^{2}}\int_{x}^{1}\mathrm{d}\xi\,\frac{x}{\xi}f_{q}^{(0)}\left(\frac{x}{\xi}\right)\left\{\frac{1+\xi^{2}}{(1-\xi)_{+}}+\epsilon(1-\xi)+\delta(1-\xi)\left[\ln\frac{Q^{2}}{P_{\perp}^{2}}+\ln\frac{1}{R^{2}}\right]\right\}\,.$ (121)
As an important consistency check, we notice that this result coincides — modulo the $\epsilon$ dependent term which vanishes in four dimension — with the corresponding one obtained in the dipole picture, cf. Eq. (11) in the Letter.
As earlier mentioned, it is convenient to transform this result to impact parameter space, cf. Eq. (113). This can be done with the help of the following identities:
$\mu^{2\epsilon}\int\frac{\mathrm{d}^{2-2\epsilon}\bm{P}_{\perp}}{(2\pi)^{2-2\epsilon}}e^{-i\bm{b}_{\perp}\cdot\bm{P}_{\perp}}\frac{1}{P_{\perp}^{2}}$ $=\frac{\pi^{-1+\epsilon}}{4}\left(\frac{\bm{b}_{\perp}^{2}\mu^{2}}{2}\right)^{\epsilon}\Gamma(-\epsilon)$ (122)
$\mu^{2\epsilon}\int\frac{\mathrm{d}^{2-2\epsilon}\bm{P}_{\perp}}{(2\pi)^{2-2\epsilon}}e^{-i\bm{b}_{\perp}\cdot\bm{P}_{\perp}}\frac{1}{P_{\perp}^{2}}\ln\left(\frac{Q^{2}}{P_{\perp}^{2}}\right)$ $=\frac{1}{(4\pi)^{1-\epsilon}}\left(\frac{\bm{b}_{\perp}^{2}\mu^{2}}{4}\right)^{\epsilon}\left[\gamma_{E}+\ln\left(\frac{Q^{2}\bm{b}_{\perp}^{2}}{4}\right)-\frac{\Gamma^{\prime}(-\epsilon)}{\Gamma(-\epsilon)}\right]\Gamma(-\epsilon)$ (123)
The full NLO correction to $x\widetilde{F}_{UU}(x,\bm{b}_{\perp},Q)$ can finally be obtained by combining the Fourier transform of Eq. (121) with the virtual corrections in Eq. (114) and Eq. (115). We thus find
$x\widetilde{F}_{UU}^{(1)}(x,\bm{b}_{\perp},Q)$ $=\,\frac{\alpha_{s}C_{F}}{2\pi}\int_{x}^{1}\mathrm{d}\xi\,\frac{x}{\xi}f_{q}^{(0)}\left(\frac{x}{\xi}\right)\Bigg{\{}\left(-\frac{1}{\epsilon}+\ln\frac{\mu_{b}^{2}}{\mu^{2}}\right)\frac{1}{C_{F}}P_{qq}^{(+)}(\xi)+(1-\xi)+$
$+\delta(1-\xi)\left[-\frac{1}{2}\left(\ln\frac{Q^{2}}{\mu_{b}^{2}}\right)^{2}+\left(\frac{3}{2}-\ln\frac{1}{R^{2}}\right)\ln\frac{Q^{2}}{\mu_{b}^{2}}+\frac{1}{2}\ln^{2}\left(\frac{1}{R^{2}}\right)+\frac{3}{2}\ln\frac{1}{R^{2}}-\frac{3}{2}-\frac{2\pi^{2}}{3}\right]\Bigg{\}}\ ,$ (124)
where $\mu_{b}=c_{0}/b_{\perp}$, $c_{0}=2e^{-\gamma_{E}}$, and $P_{qq}^{(+)}(\xi)=C_{F}\frac{1+\xi^{2}}{(1-\xi)_{+}}$ is the quark splitting kernel. The second term in the curly bracket of Eq. (124), just before the $\delta(1-\xi)$ piece, is a finite piece coming from the product of the $1/\epsilon$ pole from the Fourier transform of $1/q_{\perp}^{2}$ and the $\mathcal{O}(\epsilon)$ term in Eq. (121). Under the Fourier transform, it is legitimate to identify $\mu_{b}^{2}$ with $P_{\perp}^{2}$. With this identification, it is easy to see that the Sudakov double and single logarithms in the square brackets multiplying the $\delta$–function $\delta(1-\xi)$ coincide with the corresponding terms in Eq. (12) in the Letter.
In what follows, we will explicitly check that the Sudakov contributions aforementioned are indeed consistent with the CSS evolution of the quark TMD. To that aim, we follow the standard CSS procedure: in the TMD factorisation, the quark distribution can be defined through subtraction,
$\mathcal{F}_{q}^{\rm\,(sub.)}(x,\bm{b}_{\perp},\mu_{F},\zeta_{c})=\mathcal{F}_{q}^{\rm\,(unsub.)}(x,\bm{b}_{\perp})\sqrt{\frac{S^{\bar{n},v}(\bm{b}_{\perp})}{S^{n,\bar{n}}(\bm{b}_{\perp})S^{n,v}(\bm{b}_{\perp})}}\ .$ (125)
The un-subtracted TMD quark distribution is defined as (in momentum space)
$\mathcal{F}_{q}^{\rm\,(unsub.)}(x,\bm{P}_{\perp})$ $=$ $\frac{1}{2}\int\frac{\mathrm{d}\xi^{-}\mathrm{d}^{2}\bm{\xi}_{\perp}}{(2\pi)^{3}}e^{-ix\xi^{-}P^{+}+i\bm{\xi}_{\perp}\cdot\bm{P}_{\perp}}\left\langle P\left|\overline{\psi}(\xi){\cal L}_{n}^{\dagger}(\xi)\gamma^{+}{\cal L}_{n}(0)\psi(0)\right|P\right\rangle\ ,$ (126)
[…]
$S^{v_{1},v_{2}}(\bm{b}_{\perp})={\langle 0|{\cal L}_{v_{2}}^{\dagger}(\bm{b}_{\perp}){\cal L}_{v_{1}}^{\dagger}(\bm{b}_{\perp}){\cal L}_{v_{1}}(0){\cal L}_{v_{2}}(0)|0\rangle}\,.$ (127)
Here, $\bm{b}_{\perp}$ is the Fourier conjugate variable with respect to the transverse momentum $\bm{P}_{\perp}$ and $\mu_{F}$ is the factorization scale. And $\zeta_{c}^{2}$ is defined as $\zeta_{c}^{2}=x^{2}(2v\cdot P)^{2}/v^{2}=2(xP^{+})^{2}e^{-2y_{n}}$ with $y_{n}$ being the rapidity cut-off in the Collins-11 scheme. The second factor represents the soft factor subtraction with $n$ and $\bar{n}$ as the light-front vectors $n=(1^{-},0^{+},0_{\perp})$, $\bar{n}=(0^{-},1^{+},0_{\perp})$, whereas $v$ is an off-light-front $v=(v^{-},v^{+},0_{\perp})$ with $v^{-}\gg v^{+}$.
A standard one-loop calculation using the Collins-11 scheme yields Collins (2023); Sun and Yuan (2013)
$x\tilde{\mathcal{F}}_{q}^{\rm\,(sub.)}(x,\bm{b}_{\perp},\mu_{F})$ $=\frac{\alpha_{s}C_{F}}{2\pi}\int_{x}^{1}\mathrm{d}\xi\,\frac{x}{\xi}f_{q}\left(\frac{x}{\xi}\right)\left\{\left(-\frac{1}{\epsilon}+\ln\frac{\mu_{b}^{2}}{\mu^{2}}\right)\frac{1}{C_{F}}P_{qq}^{(+)}(\xi)+(1-\xi)\right.$
$\left.+\delta(1-\xi)\left[-\frac{1}{2}\left(\ln\frac{\zeta_{c}^{2}}{\mu_{b}^{2}}\right)^{2}+\frac{3}{2}\ln\frac{\mu_{F}^{2}}{\mu_{b}^{2}}+\frac{1}{2}\left(\ln\frac{\zeta_{c}^{2}}{\mu^{2}_{F}}\right)^{2}\right]\right\},$ (128)
where $\mu_{F}$ is the factorisation scale and $\zeta_{c}^{2}\equiv x^{2}(2v\cdot P_{N})^{2}/v^{2}=2(xP^{-}_{N})^{2}e^{-2y_{n}}$ with $y_{n}$ being the rapidity cut-off in the Collins-11 scheme ($v^{\mu}=(v^{-},v^{+},0_{\perp})$ is a slightly off-light-front 4-vector with $v^{+}\gg v^{-}$).
In order to deal with the Sudakov single log proportional to $\alpha_{s}C_{F}/(2\pi)\ln(R)$, we finally introduce the soft factor associated with the final state jet, defined according to
$\tilde{S}_{J}(\bm{b}_{\perp},\mu_{F})=\frac{\tilde{S}^{R}_{n,\bar{n}}(\bm{b}_{\perp})}{\sqrt{\tilde{S}_{n,\bar{n}}(\bm{b}_{\perp})}}\ ,$ (129)
where $R$ represents the soft gluon radiation out of the jet cone contribution. A one-loop calculation of the above soft factor associated with the jet (similar to that presented in Ref. Liu et al. (2019)) leads to the following result
$S_{J}^{(1)}(P_{\perp})=\frac{\alpha_{s}C_{F}}{2\pi^{2}}\frac{1}{P_{\perp}^{2}}\ln\frac{1}{R^{2}}\ ,$ (130)
in the transverse momentum space. In $b_{\perp}$-space, we obtain, to first order,
$\tilde{S}_{J}^{(1)}(\bm{b}_{\perp},\mu_{F})=\frac{\alpha_{s}}{2\pi}\left[-\ln\frac{1}{R^{2}}\ln\frac{\mu_{F}^{2}}{\mu_{b}^{2}}\right]\ .$ (131)
From this result, we derive the anomalous dimension at one-loop order,
$\gamma_{s}^{(1)}=-C_{F}\frac{\alpha_{s}}{2\pi}\ln\frac{1}{R^{2}}\ .$ (132)
With the above results, we can verify the TMD factorisation at this order,
$xF_{UU}(x,\bm{b}_{\perp},\mu_{F},\zeta_{c})$ $=$ $x\tilde{\mathcal{F}}_{q}^{\textrm{(sub.)}}(x,\bm{b}_{\perp},\mu_{F},\zeta_{c})\tilde{S}_{J}(\bm{b}_{\perp},\mu_{F})H_{\rm TMD}(Q,\mu_{F})\ ,$ (133)
where to the NLO accuracy of interest, $\tilde{S}_{J}\simeq 1+\tilde{S}_{J}^{(1)}$ and the hard factor takes the following form:
$H_{\rm TMD}^{(1)}$ $=$ $\frac{\alpha_{s}C_{F}}{2\pi}\left[\frac{1}{2}\ln^{2}\frac{1}{R^{2}}+\frac{3}{2}\ln\frac{1}{R^{2}}-\frac{3}{2}-\frac{2\pi^{2}}{3}\right]\ ,$ (134)
where we have chosen $\mu_{F}^{2}=\zeta_{c}^{2}=Q^{2}$ to simplify the final expression. With these particular choices, which are physically meaningful, it is easy to check that the expression for $xF_{UU}^{(1)}(x,\bm{b}_{\perp},Q^{2})$ predicted by TMD factorisation, Eq. (133), at one-loop order coincides with our NLO result for the jet cross-section, as presented in Eq. (124). For instance, the last double logarithm within the square brackets in Eq. (D) vanishes, while the first two logarithms there account for the double Sudakov logarithm and the Sudakov single log proportional to $3\alpha_{s}C_{F}/(4\pi)$ in Eq. (124). One can similarly check all the other terms in the r.h.s. of Eq. (124). Importantly, the hard factor (134) does not contain any large Sudakov logarithm of the form $\ln(Q^{2}/\mu_{b}^{2})$, which means that the CSS evolution of the quark TMD distribution correctly accounts for the Sudakov logarithms which were initially present in the NLO correction given by Eq. (124). Since the value of the coefficients of these Sudakov logarithms in Eq. (124) is a consequence of the choice of jet definition, we obtain again, but this time in the moderate $x$ regime, that our jet definition is consistent with TMD factorisation and CSS+DGLAP evolution.
To conclude this discussion, let us recall that the NLO corrections to jet production in the valence quark picture are unable to distinguish between different values of the parameter $\beta$ in the $B$ class of clustering conditions, Eq. (32). The $\beta$–dependence should however become visible at NNLO, or two loop, order. Our previous calculation using sea quarks in the dipole picture predicts that the only acceptable value (in the sense of the conformity with TMD factorisation) is $\beta=0$. If the DIS jets are clustered in the Breit frame by using spherically invariant distance measures as in Refs. Gutierrez-Reyes et al. (2018, 2019a, 2019b) (which correspond to Eq. (32) with $\beta=2$), we therefore expect violations of TMD factorisation starting with the two loops order.